### v1

# Fabian Mauthe

Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Hagmeyer, S.; Mauthe, F.; Zeiler, P. (2021): Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087.
Fabian Mauthe
Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
Goeppingen, Germany
fabian.mauthe@hs-esslingen.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Stuttgart, Germany
christopher.braun@iff.uni-stuttgart.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Stuttgart, Germany
julian.raible@iff.uni-stuttgart.de
[…]
Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
Goeppingen, Germany
peter.zeiler@hs-esslingen.de
[…]
Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
Fraunhofer Institute of Manufacturing Engineering and Automation IPA
Stuttgart, Germany
marco.huber@ieee.org
1 Prognostics and Health Management
In the realm of modern manufacturing, significant emphasis is placed on improving the reliability, performance, and service life of complex engineering systems, which has catalyzed the emergence of Prognostics and Health Management (PHM) as an essential discipline in the past two decades [1, 2]. PHM includes a range of methods dedicated to the early detection, prognosis, and mitigation of impending faults or failures in ESs, thereby facilitating sophisticated maintenance strategies and enhancing operational efficiency [3].
Central to the efficacy of PHM methods is the acquisition and analysis of degradation data, which encapsulates the evolving health condition of ESs over time [3]. Degradation data serves as a rich source of information, offering invaluable insights into the underlying degradation processes, failure modes, and performance trends of ESs [2, 4]. Through the integration of advanced sensing technologies, data analytics, and machine learning algorithms, researchers and practitioners can harness degradation data to develop prognostic models, optimize maintenance schedules, and make informed operational decisions [1].
In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview of publicly available degradation data sets is presented as well as briefly described. The authors aim to update the overview regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
2 Tasks of Diagnostics and Prognostics
Extensive and representative degradation data enables addressing a specific task—prognosis—within PHM, with the subdivision of PHM into distinct tasks. Based on [1], [5], and [4], the tasks of diagnostics and prognostics within PHM can be subdivided as follows:
  • Fault detection/anomaly detection: Detect a fault state or anomaly in an ES without considering its root cause. This results in a binary classification problem with the states fault or no fault.
  • Diagnosis: Assign one or more causes to a detected fault state.
  • Health assessment: Assess the state of degradation or the current risk of failure of an ES based on its current condition.
  • Prognosis: Prediction of the future degradation behavior or the current Remaining Useful Life (RUL).
The prognosis task poses the greatest challenge in diagnostics and prognostics within PHM [6]. This difficulty stems from the intricate nature of failure mechanisms and the multitude of factors influencing degradation progression. Therefore, sufficient degradation data is of huge importance for this task and needs to be comprehensively available.
3 Data Sets for Diagnostics and Prognostics
Table Appendix: Overview of Publicly Available Data Sets in the Appendix: Overview of Publicly Available Data Sets contains an overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically. This is the extended overview from [4], which already contains 70 data sets.
The designation of the data sets is divided into two parts, consisting of the source or platform and the title under which they were published. In addition, the overview contains the following columns:
  • Domain and Application: The associated domain from which the data set originates and the respective application within this domain.
  • Task: The corresponding task within PHM addressable with this data set.
  • URL: The respective hyperlink to the source or platform hosting the data set.
4 Description of Data Sets
The overview of publicly available degradation data sets in Table Appendix: Overview of Publicly Available Data Sets comprises a total of 79 data sets. These data sets originate from eleven different domains. The domains as well as the applications presented do not claim to be exhaustive and are based solely on the available data sets. With 17 assigned data sets, the Mechanical component domain is the most extensive, followed by the Electrical component domain with 13 assigned data sets. The two most frequently occurring applications, Bearing (ten) and Battery (six), also originate from these two domains. In total, the domains contain 36 different applications.
Except for the Building domain, which is sparsely represented in PHM, the data sets are evenly distributed for the remaining domains, with at least four and at most nine assignments per domain. With only one data set for each application, the Mechatronic system domain contains the greatest variation of applications. Moreover, there are five data sets with unknown applications and therefore, no domain can be assigned. Data sets with degradation data resulting from a simulated application are assigned to the corresponding domains but marked. With 32 assigned data sets for prognosis and 29 for diagnosis, these tasks are the most represented. For fault detection, 15 data sets are assigned, and for health assessment, 3 data sets are assigned. The Electrical component domain contains the most data sets (ten) for the task prognosis, followed by the Drive technology domain with six and the Manufacturing process as well as Process technology domains with four assigned data sets each.
Of the 17 assigned data sets in the Mechanical component domain, only two are for task prognosis. However, for the task diagnosis, the Mechanical component domain contains the most data sets with eleven assignments. Followed by the Mechatronic system domain with five and the Robotic domain with four assigned data sets. Regarding fault detection, the domains Mechanical component and Production system contain the most data sets with four assignments each. With two data sets with an Unknown domain and one data set from the Drive technology domain, the overview contains the fewest data sets for the health assessment task.
Appendix: Overview of Publicly Available Data Sets
Overview of publicly available data sets
No.
Data Set Designation
Reference to Data Origin
Domain
Application
Task URL
[…]
Table 0 – continued from previous page
No.
Data Set Designation
Reference to Data Origin
Domain
Application
Task URL
[…]
Continued on next page
[…]
1
AIDAR Lab - Air Compressor
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2016). Intelligent Condition Based Monitoring using Acoustic Signals for Air Compressors, IEEE Transactions on Reliability, vol. 65, no. 1, pp. 291-309.
Mechatronic system
Air Compressor
Diagnosis
Link
2
AIDAR Lab - Drill Bit
Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26.
Manufacturing process
Drill bit
Diagnosis
Link
3
Aramis - Data Challenge ESREL2020PSAM15
Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France).
Unknown
Simulation (time-continuous stochastic process)*
Fault detection
Link
4
Backblaze - Hard Drive Stats
Backblaze, Inc.
Drive technology
Disk Drives
Fault detection
Link
5
Bearing Data Center - Fan and Bearing
Case Western Reserve University, USA
Mechanical component
Bearing
Diagnosis
Link
6
GitHub - Predictive Maintenance using PySpark
GitHub User: linya9191
Simulation
Simulation*
Fault detection
Link
7
Kaggle - Air pressure system failures in Scania trucks
Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm
Mechatronic system
Air Pressure System
Fault detection
Link
8
Kaggle - CNC Mill Tool Wear
System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan
Manufacturing process
Mill Tool
Fault detection
Link
9
Kaggle - Condition Data with Random Recording Time
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Prognosis
Link
10
Kaggle - E-coating ultrafiltration maintenance
Tinsley Bridge Limited, Sheffield, United Kingdom
Manufacturing process
Electrophoresis Painting
Prognosis
Link
11
Kaggle - Genesis Demonstrator
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Robotic
Pick-and-Place Demonstrator
Diagnosis
Link
12
Kaggle - Microsoft Azure Predictive Maintenance
Azure AI Notebooks for Predictive Maintenance
Unknown
Unknown Machine
Fault detection
Link
13
Kaggle - One Year Industrial Component Degradation
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
Shrink-Wrapper
Fault detection
Link
14
Kaggle - Kaggle - Predictive Maintenance 1
Creators Yuan Yao and Zhao Yuqi
Production system
Log Data
Fault detection
Link
15
Kaggle - Preventive to Predictive Maintenance
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Prognosis
Link
16
Kaggle - Production Plant Data for Condition Monitoring
von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018). Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems. Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRP-CMS
Production system
Productionline
Fault detection
Link
17
Kaggle - Prognosis based on Varying Data Quality
Hagmeyer, S., Mauthe, F., Zeiler, P. (2021). Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications. International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
Process technology
Filtration
Prognosis
Link
18
Kaggle - Pump
Kaggle User: UnknownClass
Process technology
Water Pump
Fault detection
Link
19
Kaggle - Sensor Fault Detection
Schneider-Electric, France
Electrical component
Temperature Sensor
Fault detection
Link
20
Kaggle - Versatile Production System
Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
Production system
Versatile Production System
Fault detection
Link
21
Mendeley - Brushless DC motor
Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019). Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2
Drive technology
Brushless DC Motor
Prognosis
Link
22
Mendeley - Diesel Engine Faults
Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil
Drive technology
Diesel Engine
Diagnosis
Link
23
Mendeley - HUST Bearing
Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3
Mechanical component
Bearing
Diagnosis
Link
24
MFPT - Condition Based Maintenance Fault
Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems
Mechanical component
Bearing
Diagnosis
Link
25
NASA - Accelerated Battery Life Testing
Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F. (2023). Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks. Proccedings of the Annual Conference of the PHM Society, 15(1). https://doi.org/10.36001/phmconf.2023.v15i1.3463
Electrical component
Battery
Prognosis
Link
26
NASA - Bearing Data Set
J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. "Bearing Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Mechanical component
Bearing
Prognosis
Link
27
NASA - Capacitor Electrical Stress
J. Renwick, C. Kulkarni, and J Celaya "Capacitor Electrical Stress Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Electrolytic Capacitors
Prognosis
Link
28
NASA - Capacitor Electrical Stress 2
J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel "Capacitor Electrical Stress Data Set - 2", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Electrolytic Capacitors
Prognosis
Link
29
NASA - CFRP Composites Data Set
Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang "CFRP Composites Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Materials
Carbon Fibre-Reinforced Polymer Composite
Prognosis
Link
30
NASA - HIRF Battery
C. Kulkarni, E. Hogge, C. Quach and K. Goebel "HIRF Battery Data Set", NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognostic-data-repository), NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Prognosis
Link
31
NASA - IGBT
J. Celaya, Phil Wysocki, and K. Goebel "IGBT Accelerated Aging Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Transistor
Prognosis
Link
32
NASA - Li-ion Battery Aging Datasets
B. Saha and K. Goebel. "Battery Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Prognosis
Link
33
NASA - Milling Data set
A. Agogino and K. Goebel. BEST lab, UC Berkeley. "Milling Data Set ", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Manufacturing process
Milling
Progrognosis
Link
34
NASA - MOSFET Thermal Overstress Aging
J. R. Celaya, A. Saxena, S. Saha, and K. Goebel "MOSFET Thermal Overstress Aging Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Transistor
Prognosis
Link
35
NASA - Randomized Battery Usage Data Set
B. Bole, C. Kulkarni, and M. Daigle "Randomized Battery Usage Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Prognosis
Link
36
NASA - Small Satellite Power Simulation
C. Kulkarni and A. Guarneros "Small Satellite Power Simulation Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Electrical component
Battery
Diagnosis
Link
37
NASA - Turbofan engine degradation simulation data set
A. Saxena and K. Goebel. "Turbofan Engine Degradation Simulation Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
Aircraft Engine*
Prognosis
Link
38
NASA - Turbofan engine degradation simulation data set 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). "Aircraft Engine Run-to-Failure Dataset under real flight conditions", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
Drive technology
Aircraft Engine*
Prognosis
Link
39
NIST - Robot Arm Position Accuracy
National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division
Robotic
Industrial Robot
Diagnosis
Link
40
OpenEI - Evaluation of Building Fault
Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019)
Building
Building
Diagnosis
Link
41
OpenEI - Gearbox Fault Diagnosis
Yogesh Pandy, National Renewable Energy Laboratory
Mechanical component
Gear
Fault detection
Link
42
PHM Data Challenge 2008 - Turbofan
NASA Prognostics Center of Excellence
Drive technology
Aircraft Engine*
Prognosis
Link
43
PHM Data Challenge 2009 - Gearbox Fault Detection
PHM Society, Gearbox fault detection data set, 2010
Mechanical component
Gear System (Gears, Bearings, Shafts)
Diagnosis
Link
44
PHM Data Challenge 2010 - CNC milling machine cutters
X. Li, B.S. Lim, J.H. Zhou, S. Huang, S.J. Phua, K.C. Shaw, and M.J. Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798
Manufacturing process
CNC Milling Machine Cutters
Prognosis
Link
45
PHM Data Challenge 2011 - Anemometer Fault Detection
Creators Unknown
Mechanical component
Anemometer
Fault detection
Link
46
PHM Data Challenge 2013 - Unknown
Creators Unknown
Unknown
Unknown
Health assessment
Link
47
PHM Data Challenge 2014 - Unknown
Creators Unknown
Unknown
Unknown
Health assessment
Link
48
PHM Data Challenge 2015 - Plant Fault Detection
Creators Unknown
Production system
Industrial Plant Monitoring
Diagnosis
Link
49
PHM Data Challenge 2016 - Semiconductor CMP
Crystec Technology Trading GmbH
Manufacturing process
Chemicalmechanical Planarization System
Prognosis
Link
50
PHM Data Challenge 2017 - Bogie Vehicle
Creators Unknown
Mechanical component
Bogie
Diagnosis
Link
51
PHM Data Challenge 2018 - Ion Mill in Waver Manufacturing
Kai Goebel (at this time at NASA)
Production system
Ion Mill Etching Tool
Prognosis
Link
52
PHM Data Challenge 2019 - Fatigue Cracks
Creators Unknown
Materials
Aluminum Structure with Dynamic Loading
Prognosis
Link
53
PHM Data Challenge 2020 Europe - Filtration System
Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412.
Process technology
Filtration
Prognosis
Link
54
PHM Data Challenge 2021 - Turbofan 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). "Aircraft Engine Run-to-Failure Dataset under real flight conditions", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2)
Drive technology
Aircraft Engine*
Prognosis
Link
55
PHM Data Challenge 2021 Europe: SCARA-robot
Swiss Centre for Electronics and Microtechnology (CSEM)
Robotic
Industrial Robot
Diagnosis
Link
56
PHM Data Challenge 2022 - Rock Drills
Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden
Mechatronic system
Rock Drills
Diagnosis
Link
57
PHM Data Challenge 2022 Europe - PCB Production line
Danilo Giordano and Martino Trevisan, (2022). "PHME Data Challenge". European conference of the prognostics and health management society. / BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en
Production system
PCB Production Line
Diagnosis
Link
58
PHM Data Challenge 2023 - Gearbox
PHM Society, Gearbox data set, 2023
Mechanical component
Gear
Diagnosis
Link
59
PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
Japan Aerospace Exploration Agency (JAXA)
Drive technology
Propulsion System*
Health assessment
Link
60
PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
FEMTO-ST Institute, Besançon, France
Mechanical component
Bearing
Prognosis
Link
61
PHM IEEE Data Challenge 2014 - Fuel Cell
FCLAB Research Federation (FR CNRS 3539, France)
Electrical component
Fuel Cell
Prognosis
Link
62
PHM IEEE Data Challenge 2023 - Planetary Gearbox
Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC)
Mechanical component
Planetary Gear
Diagnosis
Link
63
PHM Society - Electromechanical Ball Screw Drive
GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy, under Award Number DE-AR0001290
Mechatronic system
Electromechanical Ball Screw Drive*
Diagnosis
Link
64
ResearchGate - Driveline Unbalanced Shaft
Giacomo Barbieri and David Sanchez-Londono and Laura Cattaneo and Luca Fumagalli and David Romero
Mechanical component
Unbalanced shaft (driveline)
Diagnosis
Link
65
SDOL - Diagnostics 101 bearing data
Kim, S., An, D., Choi, J-H. (2020). Diagnostics 101: A Tutorial for Fault Diagnostics of Rolling Element Bearing Using Envelope Analysis in MATLAB. Applied Sciences. 10(20):7302. doi:10.3390/app10207302
Mechanical component
Bearing
Fault detection
Link
66
SDOL - HS Gear
Creators Unknown
Mechanical component
Bearing
Fault detection
Link
67
SDOL - KAU Gear
Creators Unknown
Mechanical component
Bearing
Diagnosis
Link
68
Toyota Research Institute - Battery Cycle Life
Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-019-0356-8
Electrical component
Battery
Prognosis
Link
69
UCI - AI4I 2020 Predictive Maintenance
Matzka, S. (2020). “Explainable Artificial Intelligence for Predictive Maintenance Applications," Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp. 69-74, doi: 10.1109/AI4I49448.2020.00023 .
Production system
Manufacturing Machine
Diagnosis
Link
70
UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University
Drive technology
Naval Propulsion Plant*
Prognosis
Link
71
UCI - Condition monitoring of hydraulic systems Data Set
ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany
Mechatronic system
Hydraulic System
Prognosis
Link
72
UCI - Mechanical Analysis
F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy
Mechatronic system
Electromechanical Devices
Diagnosis
Link
73
UCI - Robot Execution Failures Data Set
Universidade Nova de Lisboa, Monte da Caparica, Portugal
Robotic
Industrial Robot
Diagnosis
Link
74
UCI - Steel Plates Faults
Semeion Research Center of Sciences of Communication - Rome, Italy
Materials
Steel Plates Faults
Diagnosis
Link
75
Uni Lulea - Vibration Signals Wind Turbines
Luleå University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements)
Mechanical component
Bearing (Wind Turbines)
Fault detection
Link
76
Uni Paderborn KAt - Bearing Damage
University Paderborn Kat
Mechanical component
Bearing
Diagnosis
Link
77
Virkler
Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153.
Materials
Aluminum Plate
Prognosis
Link
78
WASEDA - Fault Detection and Classification
De Bruijn, B., Nguyen, T. A., Bucur, D., & Tei, K. (2016). Benchmark datasets for fault detection and classification in sensor data. In A. Ahrens, O. Postolache, & C. Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Confererence on Sensor Networks, pp. 185-195. . SciTePress. https://doi.org/10.5220/0005637901850195
Electrical component
Sensors
Diagnosis
Link
79
Zenodo - Predictive maintenance dataset
Huawei Munich Research Center
Mechatronic system
Elevator
Prognosis
Link

### v2

@@
 Fabian Mauthe
-Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Hagmeyer, S.; Mauthe, F.; Zeiler, P.
-(2021): Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications.
-International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087.
+Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Fabian Mauthe, Luca Steinmann, and Peter Zeiler.
+Overview and analysis of publicly available degradation data sets for tasks within prognostics and health management.
+In Eirik Bjorheim Abrahamsen, Terje Aven, Frederic Bouder, Roger Flage, and Marja Yloenen, editors, 35th European Safety and Reliability Conference.
+(Submitted).
+Research Publishing, 2025.
 Fabian Mauthe
@@
 Through the integration of advanced sensing technologies, data analytics, and machine learning algorithms, researchers and practitioners can harness degradation data to develop prognostic models, optimize maintenance schedules, and make informed operational decisions [1].
-In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview of publicly available degradation data sets is presented as well as briefly described.
+In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview from [5] of publicly available degradation data sets is presented.
 The authors aim to update the overview regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
@@
   • Prognosis: Prediction of the future degradation behavior or the current Remaining Useful Life (RUL).
-The prognosis task poses the greatest challenge in diagnostics and prognostics within PHM [6].
-This difficulty stems from the intricate nature of failure mechanisms and the multitude of factors influencing degradation progression.
-Therefore, sufficient degradation data is of huge importance for this task and needs to be comprehensively available.
 3 Data Sets for Diagnostics and Prognostics
 Table Appendix: Overview of Publicly Available Data Sets in the Appendix: Overview of Publicly Available Data Sets contains an overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically.
-This is the extended overview from [4], which already contains 70 data sets.
+This is the extended overview from [4], presented in [5].
 The designation of the data sets is divided into two parts, consisting of the source or platform and the title under which they were published.
 In addition, the overview contains the following columns:
-  • Domain and Application: The associated domain from which the data set originates and the respective application within this domain.
+  • Domain, Application, and Signals: The associated domain from which the data set originates and the respective application within this domain.
+In addition, the signals present in the respective data set are listed.
   • Task: The corresponding task within PHM addressable with this data set.
   • URL: The respective hyperlink to the source or platform hosting the data set.
-4 Description of Data Sets
-The overview of publicly available degradation data sets in Table Appendix: Overview of Publicly Available Data Sets comprises a total of 79 data sets.
-These data sets originate from eleven different domains.
-The domains as well as the applications presented do not claim to be exhaustive and are based solely on the available data sets.
-With 17 assigned data sets, the Mechanical component domain is the most extensive, followed by the Electrical component domain with 13 assigned data sets.
-The two most frequently occurring applications, Bearing (ten) and Battery (six), also originate from these two domains.
-In total, the domains contain 36 different applications.
-Except for the Building domain, which is sparsely represented in PHM, the data sets are evenly distributed for the remaining domains, with at least four and at most nine assignments per domain.
-With only one data set for each application, the Mechatronic system domain contains the greatest variation of applications.
-Moreover, there are five data sets with unknown applications and therefore, no domain can be assigned.
-Data sets with degradation data resulting from a simulated application are assigned to the corresponding domains but marked.
-With 32 assigned data sets for prognosis and 29 for diagnosis, these tasks are the most represented.
-For fault detection, 15 data sets are assigned, and for health assessment, 3 data sets are assigned.
-The Electrical component domain contains the most data sets (ten) for the task prognosis, followed by the Drive technology domain with six and the Manufacturing process as well as Process technology domains with four assigned data sets each.
-Of the 17 assigned data sets in the Mechanical component domain, only two are for task prognosis.
-However, for the task diagnosis, the Mechanical component domain contains the most data sets with eleven assignments.
-Followed by the Mechatronic system domain with five and the Robotic domain with four assigned data sets.
-Regarding fault detection, the domains Mechanical component and Production system contain the most data sets with four assignments each.
-With two data sets with an Unknown domain and one data set from the Drive technology domain, the overview contains the fewest data sets for the health assessment task.
+The detailed PHM-specific analysis of the data sets can be found in [5].
 Appendix: Overview of Publicly Available Data Sets
-Overview of publicly available data sets
-No.
-Data Set Designation
-Reference to Data Origin
-Domain
-Application
-Task URL
+{xltabular}
+|p0.6cm|X|p4.1cm|p2.3cm|p1.0cm| Overview of publicly available data sets
+No. Data Set Designation
+Reference to Data Origin Domain
+Application and Signal Task URL
 […]
 Table 0 – continued from previous page
-No.
-Data Set Designation
-Reference to Data Origin
-Domain
-Application
-Task URL
+No. Data Set Designation
+Reference to Data Origin Domain
+Application and Signal Task URL
 […]
@@
 […]
-1
-AIDAR Lab - Air Compressor
+1 4TU - Lifecycle ageing tests on commercial 18650 Li ion cell
+Trad, Khiem (2021): Lifecycle ageing tests on commercial 18650 Li ion cell @ 25°C and 45°C. Version 1. 4TU.ResearchData. dataset. https://doi.org/10.4121/13739296.v1 Electrical component
+Battery
+Signals: current, temperature, voltage Prognosis Link
+2 4TU - Motor Current and Vibration Monitoring Dataset
+Bruinsma, Sietze; Geertsma, Rinze; Loendersloot, Richard; Tinga, Tiedo (2024): Motor Current and Vibration Monitoring Dataset for various Faults in an E-motor-driven Centrifugal Pump.
+doi: 10.4121/2b61183ec14f-4131-829b-cc4822c369d0.v3 Process technology
+Centrifugal pumps
+Signals: current, vibration, voltage Diagnosis Link
+3 AIDAR Lab - Air Compressor
 Nishchal K.
@@
 Air Compressor
-Diagnosis
-Link
-2
-AIDAR Lab - Drill Bit
-Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26.
-Manufacturing process
+Signals: acoustic emission Diagnosis Link
+4 AIDAR Lab - Drill Bit
+Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26. Manufacturing process
 Drill bit
-Diagnosis
-Link
-3
-Aramis - Data Challenge ESREL2020PSAM15
+Signals: vibration Diagnosis Link
+5 Aramis - Data Challenge ESREL2020PSAM15
 Francesco Cannarile (Aramis Srl, Italy), Michele Compare (Aramis Srl, Italy and Politecnico di Milano, Italy), Piero Baraldi (Politecnico di Milano, Italy), Zhe Yang (Politecnico di Milano, Italy), Enrico Zio (Politecnico di Milano, Italy and MINES ParisTech, France).
@@
 Simulation (time-continuous stochastic process)*
-Fault detection
-Link
-4
-Backblaze - Hard Drive Stats
-Backblaze, Inc.
-Drive technology
+Signals: vibration Fault detection Link
+6 Backblaze - Hard Drive Stats
+Backblaze, Inc. Drive technology
 Disk Drives
-Fault detection
-Link
-5
-Bearing Data Center - Fan and Bearing
-Case Western Reserve University, USA
+Signals: temperature Fault detection Link
+7 Bearing Data Center - Fan and Bearing
+Case Western Reserve University, USA Mechanical component
+Bearing
+Signals: acoustic emission Diagnosis Link
+8 Calce - Battery Data Repository
+CALCE Center for Advanced Life Cycle Engineering, University of Maryland, website: https://calce.umd.edu/battery-data Electrical component
+Battery
+Signals: current, temperature, voltage Prognosis Link
+9 GitHub - Predictive Maintenance using PySpark
+GitHub User: linya9191 Simulation
+Simulation*
+Signals: unknown Fault detection Link
+10 GitHub - XJTU-SY Bearing Datasets
+Biao Wang, Yaguo Lei, Naipeng Li, Ningbo Li (2020).
+A Hybrid Prognostics Approach for Estimating Remaining Useful Life of Rolling Element Bearings, IEEE Transactions on Reliability, vol.
+69, no.
+1, pp.
+401-412.
+DOI: 10.1109/TR.2018.2882682.
 Mechanical component
 Bearing
-Diagnosis
-Link
-6
-GitHub - Predictive Maintenance using PySpark
-GitHub User: linya9191
-Simulation
-Simulation*
-Fault detection
-Link
-7
-Kaggle - Air pressure system failures in Scania trucks
-Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm
-Mechatronic system
+Signals: vibration Prognosis Link
+11 Kaggle - Air pressure system failures in Scania trucks
+Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm Mechatronic system
 Air Pressure System
-Fault detection
-Link
-8
-Kaggle - CNC Mill Tool Wear
-System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan
-Manufacturing process
+Signals: anonymized Fault detection Link
+12 Kaggle - Bearings with Varying Degradation Behaviors
+Mauthe, F.; Hagmeyer, S.; Zeiler, P.
+(2025).
+Holistic simulation model of the temporal degradation of rolling bearings.
+In E.
+B.
+Abrahamsen, T.
+Aven, F.
+Bouder, R.
+Flage, and M.
+Yloenen (Eds.), Proceedings of the 35nd European Safety and Reliability Conference (Submitted).
+Research Publishing.
+Mechanical component
+Bearing (simulated)
+Signals: vibration Prognosis Link
+13 Kaggle - CNC Mill Tool Wear
+System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan Manufacturing process
 Mill Tool
-Fault detection
-Link
-9
-Kaggle - Condition Data with Random Recording Time
+Signals: acceleration, current, feedrate, position, speed, voltage Fault detection Link
+14 Kaggle - Condition Data with Random Recording Time
 Hagmeyer, S., Mauthe, F., Zeiler, P.
@@
 Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications.
-International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
-Process technology
+International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
 Filtration
-Prognosis
-Link
-10
-Kaggle - E-coating ultrafiltration maintenance
-Tinsley Bridge Limited, Sheffield, United Kingdom
-Manufacturing process
+Signals: flow rate, pressure Prognosis Link
+15 Kaggle - E-coating ultrafiltration maintenance
+Tinsley Bridge Limited, Sheffield, United Kingdom Manufacturing process
 Electrophoresis Painting
-Prognosis
-Link
-11
-Kaggle - Genesis Demonstrator
-Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
-Robotic
+Signals: flow rate, pressure, temperature Prognosis Link
+16 Kaggle - Genesis Demonstrator
+Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Robotic
 Pick-and-Place Demonstrator
-Diagnosis
-Link
-12
-Kaggle - Microsoft Azure Predictive Maintenance
-Azure AI Notebooks for Predictive Maintenance
-Unknown
+Signals: acceleration, current, force, speed Diagnosis Link
+17 Kaggle - Microsoft Azure Predictive Maintenance
+Azure AI Notebooks for Predictive Maintenance Unknown
 Unknown Machine
-Fault detection
-Link
-13
-Kaggle - One Year Industrial Component Degradation
-Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
-Production system
+Signals: pressure, speed, vibration, voltage Prognosis Link
+18 Kaggle - One Year Industrial Component Degradation
+Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
 Shrink-Wrapper
-Fault detection
-Link
-14
-Kaggle - Kaggle - Predictive Maintenance 1
-Creators Yuan Yao and Zhao Yuqi
-Production system
+Signals: position, speed, torque Fault detection Link
+19 Kaggle - Predictive Maintenance 1
+Creators Yuan Yao and Zhao Yuqi Production system
 Log Data
-Fault detection
-Link
-15
-Kaggle - Preventive to Predictive Maintenance
+Signals: unknown Fault detection Link
+20 Kaggle - Preventive to Predictive Maintenance
 Hagmeyer, S., Mauthe, F., Zeiler, P.
@@
 Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications.
-International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
-Process technology
+International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
 Filtration
-Prognosis
-Link
-16
-Kaggle - Production Plant Data for Condition Monitoring
+Signals: flow rate, pressure Prognosis Link
+21 Kaggle - Production Plant Data for Condition Monitoring
 von Birgelen, Alexander; Buratti, Davide; Mager, Jens; Niggemann, Oliver (2018).
 Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems.
-Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRP-CMS
-Production system
+Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRPCMS Production system
 Productionline
-Fault detection
-Link
-17
-Kaggle - Prognosis based on Varying Data Quality
+Signals: anonymized Fault detection Link
+22 Kaggle - Prognosis based on Varying Data Quality
 Hagmeyer, S., Mauthe, F., Zeiler, P.
@@
 Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications.
-International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
-Process technology
+International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
 Filtration
-Prognosis
-Link
-18
-Kaggle - Pump
-Kaggle User: UnknownClass
-Process technology
+Signals: dust feed, flow rate, pressure Prognosis Link
+23 Kaggle - Pump
+Kaggle User: UnknownClass Process technology
 Water Pump
-Fault detection
-Link
-19
-Kaggle - Sensor Fault Detection
-Schneider-Electric, France
-Electrical component
+Signals: anonymized Fault detection Link
+24 Kaggle - Sensor Fault Detection
+Schneider-Electric, France Electrical component
 Temperature Sensor
-Fault detection
-Link
-20
-Kaggle - Versatile Production System
-Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
-Production system
+Signals: temperature, voltage Fault detection Link
+25 Kaggle - Versatile Production System
+Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
 Versatile Production System
-Fault detection
-Link
-21
-Mendeley - Brushless DC motor
+Signals: control parameters Fault detection Link
+26 Mendeley - Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)
+Lu, Jiahuan; Xiong, Rui; Tian, Jinpeng; Wang, Chenxu; Hsu, Chia-Wei; Tsou, Nien-Ti; Sun, Fengchun; Li, Ju (2021), “Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)”, Mendeley Data, V2, doi: 10.17632/kw34hhw7xg.2 Electrical component
+Battery
+Signals: current, temperature, voltage Prognosis Link
+27 Mendeley - Brushless DC motor
 Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019).
-Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2
-Drive technology
+Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2 Drive technology
 Brushless DC Motor
-Prognosis
-Link
-22
-Mendeley - Diesel Engine Faults
-Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil
-Drive technology
+Signals: current, position, temperature Prognosis Link
+28 Mendeley - Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells
+Diao, Weiping (2019), “Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells”, Mendeley Data, V1,DOI: 10.17632/c35zbmn7j8.1 Electrical component
+Battery
+Signals: capacity Prognosis Link
+29 Mendeley - Diesel Engine Faults
+Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil Drive technology
 Diesel Engine
-Diagnosis
-Link
-23
-Mendeley - HUST Bearing
-Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3
+Signals: pressure, temperature, vibration Diagnosis Link
+30 Mendeley - HUST Bearing
+Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3 Mechanical component
+Bearing
+Signals: vibration Diagnosis Link
+31 Mendeley - Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell
+Zuo, Jian (2024), “Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell”, Mendeley Data, V1, doi: 10.17632/w65jjt8v5w.1 Electrical component
+Fuel Cell
+Signals: unknown Prognosis Link
+32 Mendeley - NMC cell 2600 mAh cyclic aging data
+Burzyński, Damian; Kasprzyk, Leszek (2021), “NMC cell 2600 mAh cyclic aging data”, Mendeley Data, V1, doi: 10.17632/k6v83s2xdm.1 Electrical component
+Battery
+Signals: current, temperature Prognosis Link
+33 Mendeley - Run-to-Failure Vibration Dataset of Self-Aligning Double-Row Ball Bearings
+Gabrielli, Alberto; Battarra, Mattia; Mucchi, Emiliano; Dalpiaz, Giorgio (2024).
+Physics-based prognostics of rolling-element bearings: The equivalent damaged volume algorithm, Mechanical Systems and Signal Processing, Volume 215, 2024, 111435, doi: 10.1016/j.ymssp.2024.111435.
 Mechanical component
 Bearing
-Diagnosis
-Link
-24
-MFPT - Condition Based Maintenance Fault
-Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems
-Mechanical component
+Signals: vibration Prognosis Link
+34 MFPT - Condition Based Maintenance Fault
+Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems Mechanical component
 Bearing
-Diagnosis
-Link
-25
-NASA - Accelerated Battery Life Testing
+Signals: vibration Diagnosis Link
+35 NASA - Accelerated Battery Life Testing
 Fricke, K., Nascimento, R., Corbetta, M., Kulkarni, C., Viana, F.
@@
 Proccedings of the Annual Conference of the PHM Society, 15(1).
-https://doi.org/10.36001/phmconf.2023.v15i1.3463
-Electrical component
+doi: 10.36001/phmconf.2023.v15i1.3463 Electrical component
 Battery
-Prognosis
-Link
-26
-NASA - Bearing Data Set
-J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. "Bearing Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Mechanical component
+Signals: current, temperature, voltage Prognosis Link
+36 NASA - Bearing Data Set
+J.
+Lee, H.
+Qiu, G.
+Yu, J.
+Lin, and Rexnord Technical Services.
+IMS, University of Cincinnati.
+”Bearing Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Mechanical component
 Bearing
-Prognosis
-Link
-27
-NASA - Capacitor Electrical Stress
-J. Renwick, C. Kulkarni, and J Celaya "Capacitor Electrical Stress Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Electrical component
+Signals: vibration Prognosis Link
+37 NASA - Capacitor Electrical Stress
+J. Renwick, C. Kulkarni, and J Celaya ”Capacitor Electrical Stress Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
 Electrolytic Capacitors
-Prognosis
-Link
-28
-NASA - Capacitor Electrical Stress 2
-J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel "Capacitor Electrical Stress Data Set - 2", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Electrical component
-Electrolytic Capacitors
-Prognosis
-Link
-29
-NASA - CFRP Composites Data Set
-Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang "CFRP Composites Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Materials
+Signals: impedance, voltage Prognosis Link
+38 NASA - Capacitor Electrical Stress 2
+J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel ”Capacitor Electrical Stress Data Set - 2”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
+Electrolytic Capacitor
+Signals: capacity, resistance Prognosis Link
+39 NASA - CFRP Composites Data Set
+Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang ”CFRP Composites Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Materials
 Carbon Fibre-Reinforced Polymer Composite
-Prognosis
-Link
-30
-NASA - HIRF Battery
+Signals: operating condition, strain, vibration, x-ray Prognosis Link
+40 NASA - HIRF Battery
 C.
@@
 Quach and K.
-Goebel "HIRF Battery Data Set", NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognostic-data-repository), NASA Ames Research Center, Moffett Field, CA
-Electrical component
+Goebel ”HIRF Battery Data Set”, NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognosticdata-repository), NASA Ames Research Center, Moffett Field, CA Electrical component
 Battery
-Prognosis
-Link
-31
-NASA - IGBT
-J. Celaya, Phil Wysocki, and K. Goebel "IGBT Accelerated Aging Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Electrical component
+Signals: current, temperature, voltage Prognosis Link
+41 NASA - IGBT
+J. Celaya, Phil Wysocki, and K. Goebel ”IGBT Accelerated Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
 Transistor
-Prognosis
-Link
-32
-NASA - Li-ion Battery Aging Datasets
-B. Saha and K. Goebel. "Battery Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Electrical component
+Signals: , current, temperature, voltage Prognosis Link
+42 NASA - Li-ion Battery Aging Datasets
+B. Saha and K. Goebel. ”Battery Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
 Battery
-Prognosis
-Link
-33
-NASA - Milling Data set
-A. Agogino and K. Goebel. BEST lab, UC Berkeley. "Milling Data Set ", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Manufacturing process
+Signals: current, temperature, voltage Prognosis Link
+43 NASA - Milling Data set
+A. Agogino and K. Goebel. BEST lab, UC Berkeley. ”Milling Data Set ”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Manufacturing process
 Milling
-Progrognosis
-Link
-34
-NASA - MOSFET Thermal Overstress Aging
-J. R. Celaya, A. Saxena, S. Saha, and K. Goebel "MOSFET Thermal Overstress Aging Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Electrical component
+Signals: acoustic emission, current, vibration Prognosis Link
+44 NASA - MOSFET Thermal Overstress Aging
+J. R. Celaya, A. Saxena, S. Saha, and K. Goebel ”MOSFET Thermal Overstress Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
 Transistor
-Prognosis
-Link
-35
-NASA - Randomized Battery Usage Data Set
-B. Bole, C. Kulkarni, and M. Daigle "Randomized Battery Usage Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Electrical component
+Signals: current, temperature, voltage Prognosis Link
+45 NASA - Randomized Battery Usage Data Set
+B. Bole, C. Kulkarni, and M. Daigle ”Randomized Battery Usage Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
 Battery
-Prognosis
-Link
-36
-NASA - Small Satellite Power Simulation
-C. Kulkarni and A. Guarneros "Small Satellite Power Simulation Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Electrical component
+Signals: current, temperature, voltage Prognosis Link
+46 NASA - Small Satellite Power Simulation
+C. Kulkarni and A. Guarneros ”Small Satellite Power Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
 Battery
-Diagnosis
-Link
-37
-NASA - Turbofan engine degradation simulation data set
-A. Saxena and K. Goebel. "Turbofan Engine Degradation Simulation Data Set", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Drive technology
+Signals: current, temperature, voltage Diagnosis Link
+47 NASA - Turbofan engine degradation simulation data set
+A. Saxena and K. Goebel. ”Turbofan Engine Degradation Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
 Aircraft Engine*
-Prognosis
-Link
-38
-NASA - Turbofan engine degradation simulation data set 2
+Signals: anonymized, operating condition Prognosis Link
+48 NASA - Turbofan engine degradation simulation data set 2
 M.
@@
 Fink (2021).
-"Aircraft Engine Run-to-Failure Dataset under real flight conditions", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
-Drive technology
+”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
 Aircraft Engine*
-Prognosis
-Link
-39
-NIST - Robot Arm Position Accuracy
-National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division
-Robotic
+Signals: anonymized, operating condition Prognosis Link
+49 NIST - Robot Arm Position Accuracy
+National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division Robotic
 Industrial Robot
-Diagnosis
-Link
-40
-OpenEI - Evaluation of Building Fault
-Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019)
-Building
-Building
-Diagnosis
-Link
-41
-OpenEI - Gearbox Fault Diagnosis
-Yogesh Pandy, National Renewable Energy Laboratory
-Mechanical component
+Signals: current, position, speed Diagnosis Link
+50 OpenEI - Evaluation of Building Fault
+Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019) Building
+Buildings
+Signals: pressure, speed, temperature Diagnosis Link
+51 OpenEI - Gearbox Fault Diagnosis
+Yogesh Pandy, National Renewable Energy Laboratory Mechanical component
 Gear
-Fault detection
-Link
-42
-PHM Data Challenge 2008 - Turbofan
-NASA Prognostics Center of Excellence
-Drive technology
+Signals: vibration Fault detection Link
+52 Oxford - Oxford Battery Degradation Dataset
+Birkl, C. (2017). Oxford Battery Degradation Dataset 1. University of Oxford. DOI: 10.5287/bodleian:KO2kdmYGg
+Electrical component
+Battery
+Signals: current, temperature, voltage Prognosis Link
+53 PHM Data Challenge 2008 - Turbofan
+NASA Prognostics Center of Excellence Drive technology
 Aircraft Engine*
-Prognosis
-Link
-43
-PHM Data Challenge 2009 - Gearbox Fault Detection
-PHM Society, Gearbox fault detection data set, 2010
-Mechanical component
+Signals: anonymized, operating condition Prognosis Link
+54 PHM Data Challenge 2009 - Gearbox Fault Detection
+PHM Society, Gearbox fault detection data set, 2010 Mechanical component
 Gear System (Gears, Bearings, Shafts)
-Diagnosis
-Link
-44
-PHM Data Challenge 2010 - CNC milling machine cutters
+Signals: speed, vibration Diagnosis Link
+55 PHM Data Challenge 2010 - CNC milling machine cutters
 X.
@@
 Shaw, and M.J.
-Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798
-Manufacturing process
+Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798 Manufacturing process
 CNC Milling Machine Cutters
-Prognosis
-Link
-45
-PHM Data Challenge 2011 - Anemometer Fault Detection
-Creators Unknown
-Mechanical component
+Signals: acoustic emission, force, vibration Prognosis Link
+56 PHM Data Challenge 2011 - Anemometer Fault Detection
+Creators Unknown Mechanical component
 Anemometer
-Fault detection
-Link
-46
-PHM Data Challenge 2013 - Unknown
-Creators Unknown
+Signals: direction, speed, temperature Fault detection Link
+57 PHM Data Challenge 2013 - Unknown
+Creators Unknown Unknown
 unknown
+Signals: anonymized Health assessment Link
+58 PHM Data Challenge 2014 - Unknown
+Creators Unknown Unknown
 unknown
-Health assessment
-Link
-47
-PHM Data Challenge 2014 - Unknown
-Creators Unknown
-Unknown
-Unknown
-Health assessment
-Link
-48
-PHM Data Challenge 2015 - Plant Fault Detection
-Creators Unknown
-Production system
+Signals: anonymized Health assessment Link
+59 PHM Data Challenge 2015 - Plant Fault Detection
+Creators Unknown Production system
 Industrial Plant Monitoring
-Diagnosis
-Link
-49
-PHM Data Challenge 2016 - Semiconductor CMP
-Crystec Technology Trading GmbH
-Manufacturing process
+Signals: anonymized Diagnosis Link
+60 PHM Data Challenge 2016 - Semiconductor CMP
+Crystec Technology Trading GmbH Manufacturing process
 Chemicalmechanical Planarization System
-Prognosis
-Link
-50
-PHM Data Challenge 2017 - Bogie Vehicle
-Creators Unknown
-Mechanical component
+Signals: flow rate, operating condition, pressure, speed Prognosis Link
+61 PHM Data Challenge 2017 - Bogie Vehicle
+Creators Unknown Mechanical component
 Bogie
-Diagnosis
-Link
-51
-PHM Data Challenge 2018 - Ion Mill in Waver Manufacturing
-Kai Goebel (at this time at NASA)
-Production system
+Signals: vibration Diagnosis Link
+62 PHM Data Challenge 2018 - Ion Mill in Waver Manufactoring
+Kai Goebel (at this time at NASA) Production system
 Ion Mill Etching Tool
-Prognosis
-Link
-52
-PHM Data Challenge 2019 - Fatigue Cracks
-Creators Unknown
-Materials
+Signals: current, pressure, speed Prognosis Link
+63 PHM Data Challenge 2019 - Fatigue Cracks
+Creators Unknown Materials
 Aluminum Structure with Dynamic Loading
-Prognosis
-Link
-53
-PHM Data Challenge 2020 Europe - Filtration System
-Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412.
-Process technology
+Signals: vibration Prognosis Link
+64 PHM Data Challenge 2020 Europe - Filtration System
+Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412. Process technology
 Filtration
-Prognosis
-Link
-54
-PHM Data Challenge 2021 - Turbofan 2
+Signals: flow rate, pressure Prognosis Link
+65 PHM Data Challenge 2021 - Turbofan 2
 M.
@@
 Fink (2021).
-"Aircraft Engine Run-to-Failure Dataset under real flight conditions", NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2)
-Drive technology
+”Aircraft Engine Runto-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2 ) Drive technology
 Aircraft Engine*
-Prognosis
-Link
-55
-PHM Data Challenge 2021 Europe: SCARA-robot
-Swiss Centre for Electronics and Microtechnology (CSEM)
-Robotic
+Signals: anonymized, operating condition Prognosis Link
+66 PHM Data Challenge 2021 Europe: SCARA-robot
+Swiss Centre for Electronics and Microtechnology (CSEM) Robotic
 Industrial Robot
-Diagnosis
-Link
-56
-PHM Data Challenge 2022 - Rock Drills
-Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden
-Mechatronic system
+Signals: current, duration, heat slope, humidity, position, pressure, speed, temperature Diagnosis Link
+67 PHM Data Challenge 2022 - Rock Drills
+Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden Mechatronic system
 Rock Drills
-Diagnosis
-Link
-57
-PHM Data Challenge 2022 Europe - PCB Production line
+Signals: pressure Diagnosis Link
+68 PHM Data Challenge 2022 Europe - PCB Production line
 Danilo Giordano and Martino Trevisan, (2022).
-"PHME Data Challenge".
+”PHME Data Challenge”.
 European conference of the prognostics and health management society.
-/ BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en
-Production system
+/ BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en Production system
 PCB Production Line
-Diagnosis
-Link
-58
-PHM Data Challenge 2023 - Gearbox
-PHM Society, Gearbox data set, 2023
-Mechanical component
+Signals: inspection data Diagnosis Link
+69 PHM Data Challenge 2023 - Gearbox
+PHM Society, Gearbox data set, 2023 Mechanical component
 Gear
-Diagnosis
-Link
-59
-PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
-Japan Aerospace Exploration Agency (JAXA)
-Drive technology
+Signals: speed, vibration Diagnosis Link
+70 PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
+Japan Aerospace Exploration Agency (JAXA) Drive technology
 Propulsion System*
-Health assessment
-Link
-60
-PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
-FEMTO-ST Institute, Besançon, France
-Mechanical component
+Signals: pressure Health assessment Link
+71 PHM Data Challenge 2024 - Helicopter Turbine Engines
+Prognostics and Health Management (PHM) Society Drive technology
+Helicopter Turbine Engine
+Signals: power, speed, temperature, torque Fault detection Link
+72 PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
+FEMTO-ST Institute, Besan¸con, France Mechanical component
 Bearing
-Prognosis
-Link
-61
-PHM IEEE Data Challenge 2014 - Fuel Cell
-FCLAB Research Federation (FR CNRS 3539, France)
-Electrical component
+Signals: temperature, vibration Prognosis Link
+73 PHM IEEE Data Challenge 2014 - Fuel Cell
+FCLAB Research Federation (FR CNRS 3539, France) Electrical component
 Fuel Cell
-Prognosis
-Link
-62
-PHM IEEE Data Challenge 2023 - Planetary Gearbox
-Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC)
-Mechanical component
+Signals: current, current density, flow rate, hygrometry, pressure, temperature, voltage Prognosis Link
+74 PHM IEEE Data Challenge 2023 - Planetary Gearbox
+Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC) Mechanical component
 Planetary Gear
-Diagnosis
-Link
-63
-PHM Society - Electromechanical Ball Screw Drive
-GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy, under Award Number DE-AR0001290
-Mechatronic system
+Signals: vibration Diagnosis Link
+75 PHM Society - Electromechanical Ball Screw Drive
+GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects AgencyEnergy (ARPA-E), U.S. Department of Energy, under Award Number DEAR0001290 Mechatronic system
 Electromechanical Ball Screw Drive*
-Diagnosis
-Link
-64
-ResearchGate - Driveline Unbalanced Shaft
-Giacomo Barbieri and David Sanchez-Londono and Laura Cattaneo and Luca Fumagalli and David Romero
-Mechanical component
+Signals: current, position, speed, torque Diagnosis Link
+76 ResearchGate - Driveline Unbalanced Shaft
+Giacomo Barbieri and David SanchezLondono and Laura Cattaneo and Luca Fumagalli and David Romero Mechanical component
 Unbalanced shaft (driveline)
-Diagnosis
-Link
-65
-SDOL - Diagnostics 101 bearing data
+Signals: vibration Diagnosis Link
+77 SDOL - Diagnostics 101 bearing data
 Kim, S., An, D., Choi, J-H.
@@
 10(20):7302.
-doi:10.3390/app10207302
+doi:10.3390/app10207302 Mechanical component
+Bearing
+Signals: vibration Diagnosis Link
+78 SDOL - HS Gear
+Creators Unknown Mechanical component
+Bearing
+Signals: vibration Fault detection Link
+79 SDOL - KAU Gear
+Creators Unknown Mechanical component
+Bearing
+Signals: encoder Diagnosis Link
+80 Toyota Research Institute - Battery Cycle Life
+Severson, K.A., Attia, P.M., Jin, N.
+et al.
+Data-driven prediction of battery cycle life before capacity degradation.
+Nat Energy 4, 383–391 (2019).
+https://doi.org/10.1038/s41560-0190356-8 Electrical component
+Battery
+Signals: current, temperature, voltage Prognosis Link
+81 UBFC - AMPERE Detection and diagnostics of rotor and stator faults in rotating machines
+Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06-03 Drive technology
+Squirrel cage motor (Rotor, Stator, Bearing)
+Signals: current, speed, vibration, voltage Diagnosis Link
+82 UBFC - LASPI Detection and diagnostics of gearbox faults
+Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06 .
 Mechanical component
-Bearing
-Fault detection
-Link
-66
-SDOL - HS Gear
-Creators Unknown
-Mechanical component
-Bearing
-Fault detection
-Link
-67
-SDOL - KAU Gear
-Creators Unknown
-Mechanical component
-Bearing
-Diagnosis
-Link
-68
-Toyota Research Institute - Battery Cycle Life
-Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-019-0356-8
-Electrical component
-Battery
-Prognosis
-Link
-69
-UCI - AI4I 2020 Predictive Maintenance
+Bearing + Gear
+Signals: current, vibration, voltage Diagnosis Link
+83 UBFC - METALLICADOUR Detection and diagnostics of multi-axis robot faults
+Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire G´enie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère).
+Robotic
+Multi-axes robot (cutting tool + axes drifts)
+Signals: current, force, position, torque, vibration Diagnosis Link
+84 UCI - AI4I 2020 Predictive Maintenance
 Matzka, S.
 (2020).
-“Explainable Artificial Intelligence for Predictive Maintenance Applications," Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp.
+“Explainable Artificial Intelligence for Predictive Maintenance Applications,” Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp.
 69-74, doi: 10.1109/AI4I49448.2020.00023 .
@@
 Manufacturing Machine
-Diagnosis
-Link
-70
-UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
-1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University
-Drive technology
+Signals: speed, temperature, torque, wear Diagnosis Link
+85 UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
+1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University Drive technology
 Naval Propulsion Plant*
-Prognosis
-Link
-71
-UCI - Condition monitoring of hydraulic systems Data Set
-ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany
-Mechatronic system
+Signals: flow rate, injection control, pressure, speed, temperature, torque Prognosis Link
+86 UCI - Condition monitoring of hydraulic systems Data Set
+ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany Mechatronic system
 Hydraulic System
-Prognosis
-Link
-72
-UCI - Mechanical Analysis
-F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy
-Mechatronic system
+Signals: efficiency, flow rate, power, pressure, temperature, vibration Diagnosis Link
+87 UCI - Mechanical Analysis 1990
+F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy Mechatronic system
 Electromechanical Devices
-Diagnosis
-Link
-73
-UCI - Robot Execution Failures Data Set
-Universidade Nova de Lisboa, Monte da Caparica, Portugal
-Robotic
+Signals: speed Diagnosis Link
+88 UCI - Robot Execution Failures Data Set
+Universidade Nova de Lisboa, Monte da Caparica, Portugal Robotic
 Industrial Robot
-Diagnosis
-Link
-74
-UCI - Steel Plates Faults
-Semeion Research Center of Sciences of Communication - Rome, Italy
-Materials
+Signals: force, torque Diagnosis Link
+89 UCI - Steel Plates Faults
+Semeion Research Center of Sciences of Communication - Rome, Italy Materials
 Steel Plates Faults
-Diagnosis
-Link
-75
-Uni Lulea - Vibration Signals Wind Turbines
-Luleå University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements)
-Mechanical component
+Signals: geometry, inspection data, operating condition Diagnosis Link
+90 Uni Lulea - Vibration Signals Wind Turbines
+Lulea University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements) Mechanical component
 Bearing (Wind Turbines)
-Fault detection
-Link
-76
-Uni Paderborn KAt - Bearing Damage
-University Paderborn Kat
-Mechanical component
+Signals: vibration Fault detection Link
+91 Uni Paderborn KAt - Bearing Damage
+University Paderborn Kat Mechanical component
 Bearing
-Diagnosis
-Link
-77
-Virkler
-Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153.
-Materials
+Signals: speed, temperature, torque Diagnosis Link
+92 Virkler - Fatigue Crack Propagation
+Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153. Materials
 Aluminum Plate
-Prognosis
-Link
-78
-WASEDA - Fault Detection and Classification
+Signals: crack length Prognosis Link
+93 WASEDA - Fault Detection and Classification
 De Bruijn, B., Nguyen, T.
@@
 SciTePress.
-https://doi.org/10.5220/0005637901850195
-Electrical component
+https://doi.org/10.5220/0005637901850195 Electrical component
 Sensors
-Diagnosis
-Link
-79
-Zenodo - Predictive maintenance dataset
-Huawei Munich Research Center
-Mechatronic system
+Signals: light, temperature Diagnosis Link
+94 Zenodo - Ball bearings subjected to time-varying load and speed conditions
+Aimiyekagbon, O. K. (2024). Run-tofailure data set of ball bearings subjected to time-varying load and speed conditions. Zenodo, doi: 10.5281/zenodo.10805043, University Paderborn Mechanical component
+Bearing
+Signals: temperature, vibration Prognosis Link
+95 Zenodo - Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation
+Jiangong Zhu, Yixiu Wang, Yuan Huang, R.
+Bhushan Gopaluni, Yankai Cao, Michael Heere, Martin J.
+Mühlbauer, Liuda Mereacre, Haifeng Dai, Xinhua Liu, Anatoliy Senyshyn, Xuezhe Wei, Michael Knapp, & Helmut Ehrenberg.
+(2022).
+Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation [Data set].
+Zenodo.
+https://doi.org/10.5281/zenodo.6405084 Electrical component
+Battery
+Signals: current, temperature, voltage Prognosis Link
+96 Zenodo - Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning
+Zhang, Y., Tang, Q., Zhang, Y., Wang, J., Stimming, U., & Lee, A.
+A.
+(2020).
+Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning [Data set].
+Zenodo.
+https://doi.org/10.5281/zenodo.3633835 Electrical component
+Battery
+Signals: current, temperature, voltage Prognosis Link
+97 Zenodo - Predictive maintenance dataset
+Huawei Munich Research Center Mechatronic system
 Elevator
-Prognosis
-Link
+Signals: humidity, vibration Prognosis Link
+98 Zenodo - TBSI Sunwoda Battery Dataset
+Shengyu, T. (2024). TBSI Sunwoda Battery Dataset [Data set]. Zenodo. https://doi.org/10.5281/zenodo.10715209 Electrical component
+Battery
+Signals: current, voltage Prognosis Link

### v3

@@
 Fabian Mauthe
-Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation: Fabian Mauthe, Luca Steinmann, and Peter Zeiler.
+Overview of publicly available degradation data sets for tasks within prognostics and health management thanks: Citation [6]: Fabian Mauthe, Luca Steinmann, Moritz Neu, and Peter Zeiler.
 Overview and analysis of publicly available degradation data sets for tasks within prognostics and health management.
-In Eirik Bjorheim Abrahamsen, Terje Aven, Frederic Bouder, Roger Flage, and Marja Yloenen, editors, 35th European Safety and Reliability Conference.
-(Submitted).
-Research Publishing, 2025.
+In Proceedings of the 35th European Safety and Reliability Conference and the 33rd Society for Risk Analysis Europe Conference, pages 945–952, Singapore, 2025.
+Research Publishing.
+thanks: Citation [2]: Simon Hagmeyer, Fabian Mauthe, and Peter Zeiler.
+Creation of publicly available data sets for prognostics and diagnostics addressing data scenarios relevant to industrial applications.
+International Journal of Prognostics and Health Management, 12(2):1–20, 2021.
 Fabian Mauthe
@@
 fabian.mauthe@hs-esslingen.de
-[…]
+   Christopher Braun
 Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
@@
 christopher.braun@iff.uni-stuttgart.de
-[…]
+   Julian Raible
 Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
@@
 julian.raible@iff.uni-stuttgart.de
-[…]
+   Peter Zeiler
 Institute for Technical Reliability and Prognostics IZP, Esslingen University of Applied Sciences
@@
 peter.zeiler@hs-esslingen.de
-[…]
+   Marco F. Huber
 Institute of Industrial Manufacturing and Management IFF, University of Stuttgart
@@
 Degradation data serves as a rich source of information, offering invaluable insights into the underlying degradation processes, failure modes, and performance trends of ESs [4, 2].
-Through the integration of advanced sensing technologies, data analytics, and machine learning algorithms, researchers and practitioners can harness degradation data to develop prognostic models, optimize maintenance schedules, and make informed operational decisions [1].
-In the remainder of this paper, the tasks of diagnostics and prognostics are listed, and an overview from [5] of publicly available degradation data sets is presented.
-The authors aim to update the overview regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
+The availability of appropriate degradation data remains a significant challenge in both industrial practice and academic research [10].
+Consequently, publicly available degradation data sets are of considerable value.
+They facilitate the development and empirical benchmarking of PHM methodologies and serve as a basis for demonstrating the practical applicability of the proposed approaches [5, 11, 7].
+Although publicly available degradation data sets are of significant importance, the literature addresses this subject to a limited extent.
+Overviews often consider a restricted range of sources and platforms or only specific applications, resulting in constrained summaries [8, 9].
+Analyses usually emphasize the amount of data or data quality, overlooking specific PHM aspects [3].
+As a result, the search process for suitable data sets is often very time-consuming for users of PHM methods [6].
+Therefore, the objective of this work is to provide a comprehensive overview of publicly available degradation data sets and a detailed PHM-specific analysis.
+To do this, the analysis presented in [6] is used.
+This PHM-specific analysis covers aspects that are pertinent to both users of PHM methods and of corresponding data sets.
+These aspects comprise the application represented in the data set, the originating domain of this application, the specific task the data set is designed to address, and the types of signals contained in the data set [6].
+In the remainder of this work, the tasks of diagnostics and prognostics are listed in Section 2, and an overview of publicly available degradation data sets along with a PHM-specific analysis is presented in Section 3.
+Section 4 concludes the results of this work.
+The authors aim to update the overview as well as the analysis regularly in the future, as new data sets are continually being published in the dynamic research field of PHM.
 2 Tasks of Diagnostics and Prognostics
-Extensive and representative degradation data enables addressing a specific task—prognosis—within PHM, with the subdivision of PHM into distinct tasks.
-Based on [1], [6], and [4], the tasks of diagnostics and prognostics within PHM can be subdivided as follows:
+Extensive and representative degradation data address a specific task of diagnostics and prognostics within PHM. Based on [10], [9], and [2], these tasks can be subdivided as follows:
   • Fault detection/anomaly detection: Detect a fault state or anomaly in an ES without considering its root cause. This results in a binary classification problem with the states fault or no fault.
@@
   • Health assessment: Assess the state of degradation or the current risk of failure of an ES based on its current condition.
-  • Prognosis: Prediction of the future degradation behavior or the current Remaining Useful Life (RUL).
+  • Prognosis: Prediction of future degradation behavior or the current Remaining Useful Life (RUL).
+This list of tasks can also be interpreted as a sequential framework for implementing a comprehensive PHM application, starting with fault detection and ending with prognosis [10].
+This can be illustrated by the following example of rolling bearing failure: Fault detection identifies an abnormal vibration pattern.
+Diagnosis then determines the root cause, for instance, a specific bearing component.
+The health assessment quantifies how this fault affects the bearing’s condition, while the prognosis predicts the bearing’s RUL.
+[6]
 3 Data Sets for Diagnostics and Prognostics
-Table Appendix: Overview of Publicly Available Data Sets in the Appendix: Overview of Publicly Available Data Sets contains an overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically.
-This is the extended overview from [4], presented in [5].
+3.1 Overview
+Due to regular data challenges and several research activities of the PHM community in recent years, new data sets are continuously being published.
+Therefore, the overview by Mauthe et al.
+[6] (which is already based on the overview by Hagmeyer et al.
+[2]) is updated using a similar procedure, i.e., the identical scope of platforms and sources of degradation data sets is considered [2].
+As a result, the overview is extended to 110 data sets in total.
+In accordance with [2] and [6], data sets that focus solely on process quality without including the degradation state are excluded from consideration.
+Table Appendix A: Overview of Publicly Available Data Sets in Appendix A: Overview of Publicly Available Data Sets contains the extended overview of publicly available data sets for the tasks of diagnostics and prognostics within PHM, sorted alphabetically.
+3.2 Data Set Analysis
+The detailed PHM-specific analysis of the data sets is based on the procedure presented in [6].
+The analysis is intended to shorten the time-consuming search for suitable data sets for its users.
+In particular, the focus of this analysis is on the following PHM-specific aspects [6]:
+  • Application: The application within a data set refers to the object of consideration from which the data originate.
+Thus, an application can represent a single component, such as a bearing, or an entire system, such as a production line.
+  • Domain: The domain from which the application originates. The domains associated with the previous examples would then be, for example, mechanical components or production systems.
+  • Task: The task, which is addressed by the corresponding data set. A distinction is made among the tasks described in Section 2: fault detection, diagnosis, health assessment, and prognosis.
+  • Signal: The types of signals included in the data set.
+This refers to signals that offer insight into the health state of the specific application or can be derived from it and used within PHM methods.
+Based on these PHM-specific aspects, Mauthe et al.
+[6] developed a taxonomy that is tailored to the classification of data sets in a PHM-specific manner.
+This taxonomy categorizes all data sets based on their application and domain of origin, the task addressed by the data set, and the relevant signals contained in the data set.
+The results using this taxonomy for the 110 publicly available data sets are provided in Table 1 in Appendix B: Classification of Publicly Available Data Sets and in Table 2 in Appendix C: Signals within Publicly Available Data Sets.
+The respective assignment of the PHM-specific aspects (application, domain, task, and signal) to the individual data sets can also be found in Table Appendix A: Overview of Publicly Available Data Sets in Appendix A: Overview of Publicly Available Data Sets.
+Domains and applications:
+Applying the presented taxonomy, the 110 publicly available degradation data sets are classified into 11 domains (see Table 1).
+The mechanical component domain contains the most data sets (26), followed by the electrical component domain (25) and the drive technology domain (12).
+Apart from these three largest domains, the remaining domains contain between two and eight data sets.
+Moreover, five data sets have unknown applications.
+As the respective data sets can still be used to apply PHM methods, they are taken into account and assigned to the unknown domain.
+In total, 43 different applications are included in the 11 domains, covering a variety of degradation processes.
+However, as these applications are spread across 110 data sets, the number of available data sets per application is limited.
+This is reflected in a more in-depth consideration of Table 1: The applications bearing (17 data sets) and battery (16 data sets) are considered most frequently, originating from the two main domains, mechanical component and electrical component, respectively.
+The next most frequent applications are aircraft engine, filtration, and production line, each occurring in five data sets.
+The limited number of data sets per application is further highlighted by the mechatronic system domain, which contains the most different applications (eight), with only one data set per application.
+Overall, only nine applications are covered by three or more data sets.
+Considered tasks within data sets:
+The diagnostic and prognostic tasks of the data sets [^1] Certain data sets may be assigned to several tasks.
+As these tasks build on each other, the highest-ranking task is assigned to a respective data set.
+Therefore, for each data set, only one task is assigned., as introduced in Section 2, are shown in the third column of Table 1.
+With 54 assigned data sets, the prognosis task is the most represented, followed by diagnosis with 35 entries, whereas fault detection is assigned to 18 data sets.
+As in the previous classification, the transition from diagnosis to health assessment, as well as the transition to prognosis, is often not distinct.
+Accordingly, only three data sets explicitly address the task of health assessment.
+For prognosis, the electrical component domain contributes 22 out of 54 data sets, of which 15 address battery applications.
+The mechanical component and drive technology domains follow with eight and seven prognosis data sets, respectively.
+Bearing is the second most common application for prognosis (seven data sets), while aircraft engine and filtration each occur in five prognosis data sets.
+In total, 20 distinct applications are available for prognosis.
+For diagnosis, the mechanical component domain contains the most data sets (13 out of 35), while the robotic and mechatronic system domains contain five diagnosis data sets each, and the production system domain contains three.
+The bearing application is the most frequently used for diagnosis (eight data sets), followed by articulated robot applications (four) and both gear and production line applications (three each).
+In general, 21 different applications are considered in the 35 data sets for diagnosis, indicating a wide variety of applications.
+For fault detection, the mechanical component and production system domains contribute the largest shares, with five and four data sets, respectively.
+Drive technology and manufacturing process contribute two fault detection data sets each, and the unknown domain contains two fault detection data sets without a specified application.
+Regarding applications, only bearing, milling, and production line occur in two data sets each; all remaining applications occur only once.
+Signals within data sets:
+Table 2 summarizes the 22 signals used in at least two of the given data sets.
+Note that Table 2 only lists signals that are measured directly.
+Values calculated from these, such as capacity or power, are excluded.
+Nevertheless, if only calculated signals are included in a data set, they are still listed.
+Also summarized are signals of a comparable nature, such as speed, including rotation and velocity.
+Vibration appears most frequently, occurring in 34 data sets, followed by current and temperature with 33 data sets each, and voltage with 25 data sets.
+These signals are typically measured within electrical and mechanical component domains (e.g., battery and bearing applications) and therefore are most prevalent in the analyzed data sets.
+Speed and pressure are also common, appearing in 20 and 18 data sets, respectively.
+Anonymized signals and flow rate occur in ten data sets each; however, anonymized signals permit specific observations mainly within a data set or application, limiting their utility for general analysis.
+Regarding the availability of data sets per signal, a limitation similar to applications can be observed: ten of the 22 signals occur in four or fewer data sets.
+Across tasks, signal occurrences show characteristic differences.
+For prognosis, current (22) and temperature (21) are most prevalent, followed by voltage (19) and vibration (14).
+For diagnosis, vibration dominates (16), followed by speed (11), current (nine), and temperature (eight), indicating that these signals are also relevant beyond purely electrical component data sets.
+4 Conclusion
+This work updates an overview of publicly available degradation data sets for tasks of diagnostics and prognostics within PHM and structures them using a dedicated taxonomy.
+The classification makes the current coverage across domains, applications, and tasks explicit and highlights where publicly available data are concentrated versus where only sparse coverage exists.
+Overall, the results underline that benchmarking is well supported for a few frequently studied use cases and signal modalities, while broader generalization and task-specific evaluations remain limited by uneven coverage.
+Maintaining and extending this overview will help track how coverage evolves and where additional publicly available data would have the greatest impact.
+Appendix A: Overview of Publicly Available Data Sets
 The designation of the data sets is divided into two parts, consisting of the source or platform and the title under which they were published.
@@
   • URL: The respective hyperlink to the source or platform hosting the data set.
-The detailed PHM-specific analysis of the data sets can be found in [5].
-Appendix: Overview of Publicly Available Data Sets
-{xltabular}
-|p0.6cm|X|p4.1cm|p2.3cm|p1.0cm| Overview of publicly available data sets
+Overview of publicly available data sets
 No. Data Set Designation
-Reference to Data Origin Domain
-Application and Signal Task URL
+Reference to Data Origin
+Domain
+Application and Signal
+Task URL
 […]
-Table 0 – continued from previous page
 No. Data Set Designation
-Reference to Data Origin Domain
-Application and Signal Task URL
-[…]
-  Continued on next page
+Reference to Data Origin
+Domain
+Application and Signal
+Task URL
 […]
 […]
-1 4TU - Lifecycle ageing tests on commercial 18650 Li ion cell
-Trad, Khiem (2021): Lifecycle ageing tests on commercial 18650 Li ion cell @ 25°C and 45°C. Version 1. 4TU.ResearchData. dataset. https://doi.org/10.4121/13739296.v1 Electrical component
+Trad, Khiem (2021): Lifecycle ageing tests on commercial 18650 Li ion cell @ 25°C and 45°C. Version 1. 4TU.ResearchData. dataset. https://doi.org/10.4121/13739296.v1
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 2 4TU - Motor Current and Vibration Monitoring Dataset
 Bruinsma, Sietze; Geertsma, Rinze; Loendersloot, Richard; Tinga, Tiedo (2024): Motor Current and Vibration Monitoring Dataset for various Faults in an E-motor-driven Centrifugal Pump.
-doi: 10.4121/2b61183ec14f-4131-829b-cc4822c369d0.v3 Process technology
+doi: 10.4121/2b61183ec14f-4131-829b-cc4822c369d0.v3
+Process technology
 Centrifugal pumps
-Signals: current, vibration, voltage Diagnosis Link
+Signals: current, vibration, voltage
+Diagnosis Link
 3 AIDAR Lab - Air Compressor
@@
 Air Compressor
-Signals: acoustic emission Diagnosis Link
+Signals: acoustic emission
+Diagnosis Link
 4 AIDAR Lab - Drill Bit
-Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26. Manufacturing process
+Nishchal K. Verma, R. K. Sevakula, S. Dixit and Salour A. (2015). Data Driven Approach for Drill Bit Monitoring, Reliability Digest, pp. 19-26.
+Manufacturing process
 Drill bit
-Signals: vibration Diagnosis Link
+Signals: vibration
+Diagnosis Link
 5 Aramis - Data Challenge ESREL2020PSAM15
@@
 Simulation (time-continuous stochastic process)*
-Signals: vibration Fault detection Link
+Signals: vibration
+Fault detection Link
 6 Backblaze - Hard Drive Stats
-Backblaze, Inc. Drive technology
+Backblaze, Inc.
+Drive technology
 Disk Drives
-Signals: temperature Fault detection Link
+Signals: temperature
+Fault detection Link
 7 Bearing Data Center - Fan and Bearing
-Case Western Reserve University, USA Mechanical component
+Case Western Reserve University, USA
+Mechanical component
 Bearing
-Signals: acoustic emission Diagnosis Link
+Signals: acoustic emission
+Diagnosis Link
 8 Calce - Battery Data Repository
-CALCE Center for Advanced Life Cycle Engineering, University of Maryland, website: https://calce.umd.edu/battery-data Electrical component
+CALCE Center for Advanced Life Cycle Engineering, University of Maryland, website: https://calce.umd.edu/battery-data
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
+9 ETH Zurich - Run-to-Failure High-voltage Circuit Breaker Mechanical Test Dataset
+Hsu, C.-C. (2024). Run-to-failure high-voltage circuit breaker mechanical test dataset. ETH Zürich Research Collection. DOI: 10.3929/ethz-b-000676480
+Electrical component
+Circuit breaker (high-voltage)
+Signals: current, vibration
+Prognosis Link
+10 GitHub - Motor current milling machine
+Apurv Rajeshkumar Darji (Developer), Dr. Mustafa Demetgül (Academic Supervisor)
+Manufacturing process
+Milling machine (motor current)
+Signals: current
+Fault detection Link
 11 GitHub - Predictive Maintenance using PySpark
-GitHub User: linya9191 Simulation
+GitHub User: linya9191
+Simulation
 Simulation*
-Signals: unknown Fault detection Link
+Signals: unknown
+Fault detection Link
 12 GitHub - XJTU-SY Bearing Datasets
@@
 Bearing
-Signals: vibration Prognosis Link
+Signals: vibration
+Prognosis Link
+13 Harvard Dataverse - Wind Turbine Main Bearing Fatigue Life Prediction
+Yucesan, Y. (2019). Wind Turbine Main Bearing Fatigue Life Prediction with PINN [Data set]. Harvard Dataverse. https://doi.org/10.7910/DVN/ENNXLZ
+Mechanical component
+Bearing
+Signals: speed, temperature
+Prognosis Link
 14 Kaggle - Air pressure system failures in Scania trucks
-Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm Mechatronic system
+Tony Lindgren and Jonas Biteus, Scania CV AB - Stockholm
+Mechatronic system
 Air Pressure System
-Signals: anonymized Fault detection Link
+Signals: anonymized
+Fault detection Link
 15 Kaggle - Bearings with Varying Degradation Behaviors
@@
 Flage, and M.
-Yloenen (Eds.), Proceedings of the 35nd European Safety and Reliability Conference (Submitted).
+Ylönen (Eds.), Proceedings of the 35th European Safety and Reliability Conference and the 33rd Society for Risk Analysis Europe Conference.
 Research Publishing.
+DOI: 10.3850/978-981-94-3281-3_ESREL-SRA-E2025-P8028-cd
 Mechanical component
 Bearing (simulated)
-Signals: vibration Prognosis Link
+Signals: vibration
+Prognosis Link
 16 Kaggle - CNC Mill Tool Wear
-System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan Manufacturing process
+System-level Manufacturing and Automation Research Testbed (SMART) at the University of Michigan
+Manufacturing process
 Mill Tool
-Signals: acceleration, current, feedrate, position, speed, voltage Fault detection Link
+Signals: acceleration, current, feedrate, position, speed, voltage
+Fault detection Link
 17 Kaggle - Condition Data with Random Recording Time
@@
 Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications.
-International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
+International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
+Process technology
 Filtration
-Signals: flow rate, pressure Prognosis Link
+Signals: flow rate, pressure
+Prognosis Link
 18 Kaggle - E-coating ultrafiltration maintenance
-Tinsley Bridge Limited, Sheffield, United Kingdom Manufacturing process
+Tinsley Bridge Limited, Sheffield, United Kingdom
+Manufacturing process
 Electrophoresis Painting
-Signals: flow rate, pressure, temperature Prognosis Link
+Signals: flow rate, pressure, temperature
+Prognosis Link
 19 Kaggle - Genesis Demonstrator
-Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Robotic
+Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
+Robotic
 Pick-and-Place Demonstrator
-Signals: acceleration, current, force, speed Diagnosis Link
+Signals: acceleration, current, force, speed
+Diagnosis Link
 20 Kaggle - Microsoft Azure Predictive Maintenance
-Azure AI Notebooks for Predictive Maintenance Unknown
+Azure AI Notebooks for Predictive Maintenance
+Unknown
 Unknown Machine
-Signals: pressure, speed, vibration, voltage Prognosis Link
+Signals: pressure, speed, vibration, voltage
+Prognosis Link
 21 Kaggle - One Year Industrial Component Degradation
-Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
+Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
+Production system
 Shrink-Wrapper
-Signals: position, speed, torque Fault detection Link
+Signals: position, speed, torque
+Fault detection Link
 22 Kaggle - Predictive Maintenance 1
-Creators Yuan Yao and Zhao Yuqi Production system
+Creators Yuan Yao and Zhao Yuqi
+Production system
 Log Data
-Signals: unknown Fault detection Link
+Signals: unknown
+Fault detection Link
 23 Kaggle - Preventive to Predictive Maintenance
@@
 Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications.
-International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
+International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
+Process technology
 Filtration
-Signals: flow rate, pressure Prognosis Link
+Signals: flow rate, pressure
+Prognosis Link
 24 Kaggle - Production Plant Data for Condition Monitoring
@@
 Self-Organizing Maps for Anomaly Localization and Predictive Maintenance in Cyber-Physical Production Systems.
-Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRPCMS Production system
+Proceedings of the CIRP Conference on Manufacturing Systems (CIRP CMS 2018) CIRPCMS
+Production system
 Production line
-Signals: anonymized Fault detection Link
+Signals: anonymized
+Fault detection Link
 25 Kaggle - Prognosis based on Varying Data Quality
@@
 Creation of Publicly Available Data Sets for Prognostics and Diagnostics Addressing Data Scenarios Relevant to Industrial Applications.
-International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087 Process technology
+International Journal of Prognostics and Health Management, Volume 12, Issue 2, DOI: 10.36001/ijphm.2021.v12i2.3087
+Process technology
 Filtration
-Signals: dust feed, flow rate, pressure Prognosis Link
+Signals: dust feed, flow rate, pressure
+Prognosis Link
 26 Kaggle - Pump
-Kaggle User: UnknownClass Process technology
+Kaggle User: UnknownClass
+Process technology
 Water Pump
-Signals: anonymized Fault detection Link
+Signals: anonymized
+Fault detection Link
 27 Kaggle - Sensor Fault Detection
-Schneider-Electric, France Electrical component
+Schneider-Electric, France
+Electrical component
 Temperature Sensor
-Signals: temperature, voltage Fault detection Link
+Signals: temperature, voltage
+Fault detection Link
+28 Kaggle - Similar System Data Set for Condition Prognosis
+Braig, M.; Zeiler, P.
+(2025): A Study on Using Transfer Learning to Utilize Information from Similar Systems for Data-Driven Condition Diagnosis and Prognosis.
+IEEE Access, Volume 13, pp.
+98485-98503, DOI: 10.1109/ACCESS.2025.3576435
+Process technology
+Filtration
+Signals: flow rate, pressure
+Prognosis Link
+29 Kaggle - Similar System Data Set for Fault Diagnosis
+Braig, M.; Zeiler, P.
+(2025): A Study on Using Transfer Learning to Utilize Information from Similar Systems for Data-Driven Condition Diagnosis and Prognosis.
+IEEE Access, Volume 13, pp.
+98485-98503, DOI: 10.1109/ACCESS.2025.3576435
+Mechanical component
+Bearing
+Signals: vibration
+Diagnosis Link
 30 Kaggle - Versatile Production System
-Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE) Production system
+Institut für industrielle Informationstechnik (inIT) der Technischen Hochschule Ostwestfalen-Lippe (IMPROVE)
+Production system
 Versatile Production System
-Signals: control parameters Fault detection Link
+Signals: control parameters
+Fault detection Link
+31 Karlsruhe Institute of Technology - Ball Screw Drive Surface Defect Dataset
+Schlagenhauf, T. (2021): Ball Screw Drive Surface Defect Dataset for Classification. Karlsruhe Institute of Technology (KIT). DOI: 10.5445/IR/1000133819
+Mechanical component
+Ball screw drive
+Signals: image
+Fault detection Link
+32 Karlsruhe Institute of Technology - Industrial Machine Tool Element Surface Defect Dataset
+Schlagenhauf, T.; Landwehr, M.; Fleischer, J. (2021): Industrial Machine Tool Element Surface Defect Dataset. Karlsruhe Institute of Technology (KIT). DOI: 10.5445/IR/1000129520
+Mechanical component
+Ball screw drive
+Signals: image
+Prognosis Link
 33 Mendeley - Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)
-Lu, Jiahuan; Xiong, Rui; Tian, Jinpeng; Wang, Chenxu; Hsu, Chia-Wei; Tsou, Nien-Ti; Sun, Fengchun; Li, Ju (2021), “Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)”, Mendeley Data, V2, doi: 10.17632/kw34hhw7xg.2 Electrical component
+Lu, Jiahuan; Xiong, Rui; Tian, Jinpeng; Wang, Chenxu; Hsu, Chia-Wei; Tsou, Nien-Ti; Sun, Fengchun; Li, Ju (2021), “Battery Degradation Dataset (Fixed Current Profiles & Arbitrary Uses Profiles)”, Mendeley Data, V2, doi: 10.17632/kw34hhw7xg.2
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 34 Mendeley - Brushless DC motor
 Mazzoleni, Mirko; Scandella, Matteo; Previdi, Fabio; Pispola, Giulio (2019).
-Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58 bdhy2df.2 Drive technology
+Data for: First endurance activity of a Brushless DC motor for aerospace applications - REPRISE project, Mendeley Data, V2, doi: 10.17632/m58bdhy2df.2
+Drive technology
 Brushless DC Motor
-Signals: current, position, temperature Prognosis Link
+Signals: current, position, temperature
+Prognosis Link
 35 Mendeley - Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells
-Diao, Weiping (2019), “Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells”, Mendeley Data, V1,DOI: 10.17632/c35zbmn7j8.1 Electrical component
+Diao, Weiping (2019), “Data for: Accelerated Cycle Life Testing and Capacity Degradation Modeling of LiCoO2-graphite Cells”, Mendeley Data, V1,DOI: 10.17632/c35zbmn7j8.1
+Electrical component
 Battery
-Signals: capacity Prognosis Link
+Signals: capacity
+Prognosis Link
 36 Mendeley - Diesel Engine Faults
-Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil Drive technology
+Denys Pestana-Viana - Federal Center of Technological Education Celso Suckow da Fonseca (CEFET-RJ), Rio de Janeiro, Brazil
+Drive technology
 Diesel Engine
-Signals: pressure, temperature, vibration Diagnosis Link
+Signals: pressure, temperature, vibration
+Diagnosis Link
 37 Mendeley - HUST Bearing
-Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3 Mechanical component
+Hong, Hoang Si; Thuan, Nguyen (2023), “HUST bearing: a practical dataset for ball bearing fault diagnosis”, Mendeley Data, V3, doi: 10.17632/cbv7jyx4p9.3
+Mechanical component
 Bearing
-Signals: vibration Diagnosis Link
+Signals: vibration
+Diagnosis Link
 38 Mendeley - Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell
-Zuo, Jian (2024), “Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell”, Mendeley Data, V1, doi: 10.17632/w65jjt8v5w.1 Electrical component
+Zuo, Jian (2024), “Long-term Dynamic Durability Test Dataset for Single Proton Exchange Membrane Fuel Cell”, Mendeley Data, V1, doi: 10.17632/w65jjt8v5w.1
+Electrical component
 Fuel Cell
-Signals: unknown Prognosis Link
+Signals: unknown
+Prognosis Link
 39 Mendeley - NMC cell 2600 mAh cyclic aging data
-Burzyński, Damian; Kasprzyk, Leszek (2021), “NMC cell 2600 mAh cyclic aging data”, Mendeley Data, V1, doi: 10.17632/k6v83s2xdm.1 Electrical component
+Burzyński, Damian; Kasprzyk, Leszek (2021), “NMC cell 2600 mAh cyclic aging data”, Mendeley Data, V1, doi: 10.17632/k6v83s2xdm.1
+Electrical component
 Battery
-Signals: current, temperature Prognosis Link
+Signals: current, temperature
+Prognosis Link
 40 Mendeley - Run-to-Failure Vibration Dataset of Self-Aligning Double-Row Ball Bearings
@@
 Bearing
-Signals: vibration Prognosis Link
+Signals: vibration
+Prognosis Link
 41 MFPT - Condition Based Maintenance Fault
-Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems Mechanical component
+Data Assembled and Prepared on behalf of MFPT by Dr Eric Bechhoefer, Chief Engineer, NRG Systems
+Mechanical component
 Bearing
-Signals: vibration Diagnosis Link
+Signals: vibration
+Diagnosis Link
 42 NASA - Accelerated Battery Life Testing
@@
 Prognosis of Li-ion Batteries Under Large Load Variations Using Hybrid Physics-Informed Neural Networks.
-Proccedings of the Annual Conference of the PHM Society, 15(1).
-doi: 10.36001/phmconf.2023.v15i1.3463 Electrical component
+Proceedings of the Annual Conference of the PHM Society, 15(1).
+doi: 10.36001/phmconf.2023.v15i1.3463
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 43 NASA - Bearing Data Set
-J.
-Lee, H.
-Qiu, G.
-Yu, J.
-Lin, and Rexnord Technical Services.
-IMS, University of Cincinnati.
-”Bearing Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Mechanical component
+J. Lee, H. Qiu, G. Yu, J. Lin, and Rexnord Technical Services. IMS, University of Cincinnati. ”Bearing Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Mechanical component
 Bearing
-Signals: vibration Prognosis Link
+Signals: vibration
+Prognosis Link
 44 NASA - Capacitor Electrical Stress
-J. Renwick, C. Kulkarni, and J Celaya ”Capacitor Electrical Stress Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
+J. Renwick, C. Kulkarni, and J Celaya ”Capacitor Electrical Stress Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Electrical component
 Electrolytic Capacitors
-Signals: impedance, voltage Prognosis Link
+Signals: impedance, voltage
+Prognosis Link
 45 NASA - Capacitor Electrical Stress 2
-J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel ”Capacitor Electrical Stress Data Set - 2”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
+J. Celaya, C. Kulkarni, G. Biswas, and K. Goebel ”Capacitor Electrical Stress Data Set - 2”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Electrical component
 Electrolytic Capacitor
-Signals: capacity, resistance Prognosis Link
+Signals: capacity, resistance
+Prognosis Link
 46 NASA - CFRP Composites Data Set
-Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang ”CFRP Composites Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Materials
+Abhinav Saxena, Kai Goebel, Cecilia C. Larrosa, and Fu-Kuo Chang ”CFRP Composites Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Materials
 Carbon Fibre-Reinforced Polymer Composite
-Signals: operating condition, strain, vibration, x-ray Prognosis Link
+Signals: operating condition, strain, vibration, x-ray
+Prognosis Link
 47 NASA - HIRF Battery
@@
 Quach and K.
-Goebel ”HIRF Battery Data Set”, NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognosticdata-repository), NASA Ames Research Center, Moffett Field, CA Electrical component
+Goebel ”HIRF Battery Data Set”, NASA Ames Prognostics Data Repository (http://ti.arc.nasa.gov/project/prognosticdata-repository), NASA Ames Research Center, Moffett Field, CA
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 48 NASA - IGBT
-J. Celaya, Phil Wysocki, and K. Goebel ”IGBT Accelerated Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
+J. Celaya, Phil Wysocki, and K. Goebel ”IGBT Accelerated Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Electrical component
 Transistor
-Signals: , current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 49 NASA - Li-ion Battery Aging Datasets
-B. Saha and K. Goebel. ”Battery Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
+B. Saha and K. Goebel. ”Battery Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 50 NASA - Milling Data set
-A. Agogino and K. Goebel. BEST lab, UC Berkeley. ”Milling Data Set ”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Manufacturing process
+A. Agogino and K. Goebel. BEST lab, UC Berkeley. ”Milling Data Set ”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Manufacturing process
 Milling
-Signals: acoustic emission, current, vibration Prognosis Link
+Signals: acoustic emission, current, vibration
+Prognosis Link
 51 NASA - MOSFET Thermal Overstress Aging
-J. R. Celaya, A. Saxena, S. Saha, and K. Goebel ”MOSFET Thermal Overstress Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
+J. R. Celaya, A. Saxena, S. Saha, and K. Goebel ”MOSFET Thermal Overstress Aging Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Electrical component
 Transistor
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 52 NASA - Randomized Battery Usage Data Set
-B. Bole, C. Kulkarni, and M. Daigle ”Randomized Battery Usage Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
+B. Bole, C. Kulkarni, and M. Daigle ”Randomized Battery Usage Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 53 NASA - Small Satellite Power Simulation
-C. Kulkarni and A. Guarneros ”Small Satellite Power Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Electrical component
+C. Kulkarni and A. Guarneros ”Small Satellite Power Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Electrical component
 Battery
-Signals: current, temperature, voltage Diagnosis Link
+Signals: current, temperature, voltage
+Diagnosis Link
 54 NASA - Turbofan engine degradation simulation data set
-A. Saxena and K. Goebel. ”Turbofan Engine Degradation Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
+A. Saxena and K. Goebel. ”Turbofan Engine Degradation Simulation Data Set”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Drive technology
 Aircraft Engine*
-Signals: anonymized, operating condition Prognosis Link
+Signals: anonymized, operating condition
+Prognosis Link
 55 NASA - Turbofan engine degradation simulation data set 2
@@
 Fink (2021).
-”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA Drive technology
+”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA
+Drive technology
 Aircraft Engine*
-Signals: anonymized, operating condition Prognosis Link
+Signals: anonymized, operating condition
+Prognosis Link
 56 NIST - Robot Arm Position Accuracy
-National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division Robotic
+National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division
+Robotic
 Industrial Robot
-Signals: current, position, speed Diagnosis Link
+Signals: current, position, speed
+Diagnosis Link
 57 OpenEI - Evaluation of Building Fault
-Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019) Building
+Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019)
+Building
 Buildings
-Signals: pressure, speed, temperature Diagnosis Link
+Signals: pressure, speed, temperature
+Diagnosis Link
 58 OpenEI - Gearbox Fault Diagnosis
-Yogesh Pandy, National Renewable Energy Laboratory Mechanical component
+Yogesh Pandy, National Renewable Energy Laboratory
+Mechanical component
 Gear
-Signals: vibration Fault detection Link
+Signals: vibration
+Fault detection Link
+59 OSF - Second-life lithium-ion battery aging dataset based on grid storage cycling
+Khan, M. A., & Onori, S. (2025). Second-life lithium-ion battery aging dataset based on grid storage cycling. https://doi.org/10.17605/OSF.IO/8JNR5
+Electrical component
+Battery
+Signals: current, impedance, voltage
+Prognosis Link
 60 Oxford - Oxford Battery Degradation Dataset
@@
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 61 PHM Data Challenge 2008 - Turbofan
-NASA Prognostics Center of Excellence Drive technology
+NASA Prognostics Center of Excellence
+Drive technology
 Aircraft Engine*
-Signals: anonymized, operating condition Prognosis Link
+Signals: anonymized, operating condition
+Prognosis Link
 62 PHM Data Challenge 2009 - Gearbox Fault Detection
-PHM Society, Gearbox fault detection data set, 2010 Mechanical component
+PHM Society, Gearbox fault detection data set, 2010
+Mechanical component
 Gear System (Gears, Bearings, Shafts)
-Signals: speed, vibration Diagnosis Link
+Signals: speed, vibration
+Diagnosis Link
 63 PHM Data Challenge 2010 - CNC milling machine cutters
@@
 Shaw, and M.J.
-Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798 Manufacturing process
+Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798
+Manufacturing process
 CNC Milling Machine Cutters
-Signals: acoustic emission, force, vibration Prognosis Link
+Signals: acoustic emission, force, vibration
+Prognosis Link
 64 PHM Data Challenge 2011 - Anemometer Fault Detection
-Creators Unknown Mechanical component
+Creators Unknown
+Mechanical component
 Anemometer
-Signals: direction, speed, temperature Fault detection Link
+Signals: direction, speed, temperature
+Fault detection Link
 65 PHM Data Challenge 2013 - Unknown
-Creators Unknown Unknown
+Creators Unknown
 Unknown
-Signals: anonymized Health assessment Link
+unknown
+Signals: anonymized
+Health assessment Link
 66 PHM Data Challenge 2014 - Unknown
-Creators Unknown Unknown
+Creators Unknown
 Unknown
-Signals: anonymized Health assessment Link
+unknown
+Signals: anonymized
+Health assessment Link
 67 PHM Data Challenge 2015 - Plant Fault Detection
-Creators Unknown Production system
+Creators Unknown
+Production system
 Industrial Plant Monitoring
-Signals: anonymized Diagnosis Link
+Signals: anonymized
+Diagnosis Link
 68 PHM Data Challenge 2016 - Semiconductor CMP
-Crystec Technology Trading GmbH Manufacturing process
+Crystec Technology Trading GmbH
+Manufacturing process
 Chemical mechanical planarization system
-Signals: flow rate, operating condition, pressure, speed Prognosis Link
+Signals: flow rate, operating condition, pressure, speed
+Prognosis Link
 69 PHM Data Challenge 2017 - Bogie Vehicle
-Creators Unknown Mechanical component
+Creators Unknown
+Mechanical component
 Bogie
-Signals: vibration Diagnosis Link
-62 PHM Data Challenge 2018 - Ion Mill in Waver Manufactoring
-Kai Goebel (at this time at NASA) Production system
+Signals: vibration
+Diagnosis Link
+70 PHM Data Challenge 2018 - Ion Mill in Wafer Manufacturing
+Kai Goebel (at this time at NASA)
+Production system
 Ion Mill Etching Tool
-Signals: current, pressure, speed Prognosis Link
+Signals: current, pressure, speed
+Prognosis Link
 71 PHM Data Challenge 2019 - Fatigue Cracks
-Creators Unknown Materials
+He J, Guan X, Peng T, Liu Y, Saxena A, Celaya J, Goebel K (2013).
+A multi-feature integration method for fatigue crack detection and crack length estimation in riveted lap joints using Lamb waves.
+Smart Materials and Structures.
+22(10):105007.
+Materials
 Aluminum Structure with Dynamic Loading
-Signals: vibration Prognosis Link
+Signals: vibration
+Prognosis Link
 72 PHM Data Challenge 2020 Europe - Filtration System
-Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412. Process technology
+Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412.
+Process technology
 Filtration
-Signals: flow rate, pressure Prognosis Link
+Signals: flow rate, pressure
+Prognosis Link
 73 PHM Data Challenge 2021 - Turbofan 2
@@
 Fink (2021).
-”Aircraft Engine Runto-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2 ) Drive technology
+”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2 )
+Drive technology
 Aircraft Engine*
-Signals: anonymized, operating condition Prognosis Link
+Signals: anonymized, operating condition
+Prognosis Link
 74 PHM Data Challenge 2021 Europe: SCARA-robot
-Swiss Centre for Electronics and Microtechnology (CSEM) Robotic
+Swiss Centre for Electronics and Microtechnology (CSEM)
+Robotic
 Industrial Robot
-Signals: current, duration, heat slope, humidity, position, pressure, speed, temperature Diagnosis Link
+Signals: current, duration, heat slope, humidity, position, pressure, speed, temperature
+Diagnosis Link
 75 PHM Data Challenge 2022 - Rock Drills
-Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden Mechatronic system
+Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden
+Mechatronic system
 Rock Drills
-Signals: pressure Diagnosis Link
+Signals: pressure
+Diagnosis Link
 76 PHM Data Challenge 2022 Europe - PCB Production line
@@
 European conference of the prognostics and health management society.
-/ BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en Production system
+/ BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en
+Production system
 PCB Production Line
-Signals: inspection data Diagnosis Link
+Signals: inspection data
+Diagnosis Link
 77 PHM Data Challenge 2023 - Gearbox
-PHM Society, Gearbox data set, 2023 Mechanical component
+PHM Society, Gearbox data set, 2023
+Mechanical component
 Gear
-Signals: speed, vibration Diagnosis Link
+Signals: speed, vibration
+Diagnosis Link
 78 PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
-Japan Aerospace Exploration Agency (JAXA) Drive technology
+Japan Aerospace Exploration Agency (JAXA)
+Drive technology
 Propulsion System*
-Signals: pressure Health assessment Link
+Signals: pressure
+Health assessment Link
 79 PHM Data Challenge 2024 - Helicopter Turbine Engines
-Prognostics and Health Management (PHM) Society Drive technology
+Prognostics and Health Management (PHM) Society
+Drive technology
 Helicopter Turbine Engine
-Signals: power, speed, temperature, torque Fault detection Link
+Signals: power, speed, temperature, torque
+Fault detection Link
+80 PHM Data Challenge 2025 - Aircraft Engines
+Creators Unknown, data generation: NASA AGTF30 Simulation – MATLAB Executable Steady-State Solver and Linearization Tool for the AGTF30 Engine Simulation (MEXLIN-AGTF30), https://software.nasa.gov/software/LEW-20688-1.
+See also https://software.nasa.gov/software/LEW-19717-1.
+MATLAB T-MATS Toolbox – https://ntrs.nasa.gov/api/citations/20180002976/downloads/20180002976.pdf
+Drive technology
+Aircraft Engine*
+Signals: operation conditions, pressure
+Prognosis Link
+81 PHM Data Challenge 2025 Asia Pacific - Predicting Cutter Flank Wear
+PHM Asia Pacific
+Manufacturing process
+Turning machine cutter flank wear
+Signals: acoustic emission, vibration
+Prognosis Link
+82 PHM Data Challenge 2026 Europe - Subway ticket validation door
+Soualhi, M. (2026): PHME Data Challenge. European conference of the prognostics and health management society
+Mechatronic system
+Validation door
+Signals: current, position, voltage
+Prognosis Link
 83 PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
-FEMTO-ST Institute, Besan¸con, France Mechanical component
+FEMTO-ST Institute, Besançon, France
+Mechanical component
 Bearing
-Signals: temperature, vibration Prognosis Link
+Signals: temperature, vibration
+Prognosis Link
 84 PHM IEEE Data Challenge 2014 - Fuel Cell
-FCLAB Research Federation (FR CNRS 3539, France) Electrical component
+FCLAB Research Federation (FR CNRS 3539, France)
+Electrical component
 Fuel Cell
-Signals: current, current density, flow rate, hygrometry, pressure, temperature, voltage Prognosis Link
+Signals: current, current density, flow rate, hygrometry, pressure, temperature, voltage
+Prognosis Link
 85 PHM IEEE Data Challenge 2023 - Planetary Gearbox
-Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC) Mechanical component
+Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC)
+Mechanical component
 Planetary Gear
-Signals: vibration Diagnosis Link
+Signals: vibration
+Diagnosis Link
 86 PHM Society - Electromechanical Ball Screw Drive
-GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects AgencyEnergy (ARPA-E), U.S. Department of Energy, under Award Number DEAR0001290 Mechatronic system
-Electromechanical Ball Screw Drive*
-Signals: current, position, speed, torque Diagnosis Link
+GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy, under Award Number DEAR0001290
+Mechatronic system
+Ball Screw Drive* (Electromechanical)
+Signals: current, position, speed, torque
+Diagnosis Link
 87 ResearchGate - Driveline Unbalanced Shaft
-Giacomo Barbieri and David SanchezLondono and Laura Cattaneo and Luca Fumagalli and David Romero Mechanical component
+Giacomo Barbieri and David Sanchez-Londono and Laura Cattaneo and Luca Fumagalli and David Romero
+Mechanical component
 Unbalanced shaft (driveline)
-Signals: vibration Diagnosis Link
+Signals: vibration
+Diagnosis Link
 88 SDOL - Diagnostics 101 bearing data
@@
 10(20):7302.
-doi:10.3390/app10207302 Mechanical component
+doi:10.3390/app10207302
+Mechanical component
 Bearing
-Signals: vibration Diagnosis Link
+Signals: vibration
+Diagnosis Link
 89 SDOL - HS Gear
-Creators Unknown Mechanical component
+Creators Unknown
+Mechanical component
 Bearing
-Signals: vibration Fault detection Link
+Signals: vibration
+Fault detection Link
 90 SDOL - KAU Gear
-Creators Unknown Mechanical component
+Creators Unknown
+Mechanical component
 Bearing
-Signals: encoder Diagnosis Link
+Signals: encoder
+Diagnosis Link
 91 Toyota Research Institute - Battery Cycle Life
-Severson, K.A., Attia, P.M., Jin, N.
-et al.
-Data-driven prediction of battery cycle life before capacity degradation.
-Nat Energy 4, 383–391 (2019).
-https://doi.org/10.1038/s41560-0190356-8 Electrical component
+Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-0190356-8
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 92 UBFC - AMPERE Detection and diagnostics of rotor and stator faults in rotating machines
-Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06-03 Drive technology
+Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06-03
+Drive technology
 Squirrel cage motor (Rotor, Stator, Bearing)
-Signals: current, speed, vibration, voltage Diagnosis Link
+Signals: current, speed, vibration, voltage
+Diagnosis Link
 93 UBFC - LASPI Detection and diagnostics of gearbox faults
@@
 Bearing + Gear
-Signals: current, vibration, voltage Diagnosis Link
+Signals: current, vibration, voltage
+Diagnosis Link
 94 UBFC - METALLICADOUR Detection and diagnostics of multi-axis robot faults
-Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire G´enie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère).
+Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère).
 Robotic
 Multi-axes robot (cutting tool + axes drifts)
-Signals: current, force, position, torque, vibration Diagnosis Link
+Signals: current, force, position, torque, vibration
+Diagnosis Link
 95 UCI - AI4I 2020 Predictive Maintenance
@@
 Manufacturing Machine
-Signals: speed, temperature, torque, wear Diagnosis Link
+Signals: speed, temperature, torque, wear
+Diagnosis Link
 96 UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
-1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University Drive technology
+1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University
+Drive technology
 Naval Propulsion Plant*
-Signals: flow rate, injection control, pressure, speed, temperature, torque Prognosis Link
+Signals: flow rate, injection control, pressure, speed, temperature, torque
+Prognosis Link
 97 UCI - Condition monitoring of hydraulic systems Data Set
-ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany Mechatronic system
+ZeMA - Center for Mechatronics and Automation Technology Saarbrücken, Germany
+Mechatronic system
 Hydraulic System
-Signals: efficiency, flow rate, power, pressure, temperature, vibration Diagnosis Link
+Signals: efficiency, flow rate, power, pressure, temperature, vibration
+Diagnosis Link
 98 UCI - Mechanical Analysis 1990
-F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy Mechatronic system
+F. Bergadano, A. Giordana, L. Saitta, (1990). University of Torino, Italy
+Mechatronic system
 Electromechanical Devices
-Signals: speed Diagnosis Link
+Signals: speed
+Diagnosis Link
 99 UCI - Robot Execution Failures Data Set
-Universidade Nova de Lisboa, Monte da Caparica, Portugal Robotic
+Universidade Nova de Lisboa, Monte da Caparica, Portugal
+Robotic
 Industrial Robot
-Signals: force, torque Diagnosis Link
+Signals: force, torque
+Diagnosis Link
 100 UCI - Steel Plates Faults
-Semeion Research Center of Sciences of Communication - Rome, Italy Materials
+Semeion Research Center of Sciences of Communication - Rome, Italy
+Materials
 Steel Plates Faults
-Signals: geometry, inspection data, operating condition Diagnosis Link
+Signals: geometry, inspection data, operating condition
+Diagnosis Link
 101 Uni Lulea - Vibration Signals Wind Turbines
-Lulea University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements) Mechanical component
+Lulea University of Technology: Martin del Campo Barraza, Sergio (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Sandin, Fredrik (Department of Computer Science, Electrical and Space Engineering, Embedded Internet Systems Lab); Strömbergsson, Daniel (Department of Engineering Sciences and Mathematics, Machine Elements)
+Mechanical component
 Bearing (Wind Turbines)
-Signals: vibration Fault detection Link
+Signals: vibration
+Fault detection Link
 102 Uni Paderborn KAt - Bearing Damage
-University Paderborn Kat Mechanical component
+University Paderborn Kat
+Mechanical component
 Bearing
-Signals: speed, temperature, torque Diagnosis Link
+Signals: speed, temperature, torque
+Diagnosis Link
 103 Virkler - Fatigue Crack Propagation
-Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153. Materials
+Virkler, D. A., Hillberry, B. M. and Goel, P. K. (1979). The Statistical Nature of Fatigue Crack Propagation. Journal of Engineering Materials and Technology, vol. 101, 148–153.
+Materials
 Aluminum Plate
-Signals: crack length Prognosis Link
+Signals: crack length
+Prognosis Link
 104 WASEDA - Fault Detection and Classification
@@
 Postolache, & C.
-Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Confererence on Sensor Networks, pp.
+Benavente-Peces (Eds.), SENSORNETS 2016 - Proceedings of the 5th International Conference on Sensor Networks, pp.
 185-195.
-.
 SciTePress.
-https://doi.org/10.5220/0005637901850195 Electrical component
+https://doi.org/10.5220/0005637901850195
+Electrical component
 Sensors
-Signals: light, temperature Diagnosis Link
+Signals: light, temperature
+Diagnosis Link
 105 Zenodo - Ball bearings subjected to time-varying load and speed conditions
-Aimiyekagbon, O. K. (2024). Run-tofailure data set of ball bearings subjected to time-varying load and speed conditions. Zenodo, doi: 10.5281/zenodo.10805043, University Paderborn Mechanical component
+Aimiyekagbon, O. K. (2024). Run-to-failure data set of ball bearings subjected to time-varying load and speed conditions. Zenodo, doi: 10.5281/zenodo.10805043, University Paderborn
+Mechanical component
 Bearing
-Signals: temperature, vibration Prognosis Link
+Signals: temperature, vibration
+Prognosis Link
 106 Zenodo - Data-driven capacity estimation of commercial lithium-ion batteries from voltage relaxation
@@
 Zenodo.
-https://doi.org/10.5281/zenodo.6405084 Electrical component
+https://doi.org/10.5281/zenodo.6405084
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
 107 Zenodo - Identifying degradation patterns of lithium ion batteries from impedance spectroscopy using machine learning
@@
 Zenodo.
-https://doi.org/10.5281/zenodo.3633835 Electrical component
+https://doi.org/10.5281/zenodo.3633835
+Electrical component
 Battery
-Signals: current, temperature, voltage Prognosis Link
+Signals: current, temperature, voltage
+Prognosis Link
+108 Zenodo - Monitoring data railway bridge Leuven
+Kristof Maes, & Geert Lombaert. (2020). Monitoring data for railway bridge KW51 in Leuven, Belgium, before, during, and after retrofitting [Data set]. Zenodo. https://doi.org/10.5281/zenodo.3745914
+Building
+Railway bridge
+Signals: acceleration, displacement, strain
+Diagnosis Link
 109 Zenodo - Predictive maintenance dataset
-Huawei Munich Research Center Mechatronic system
+Huawei Munich Research Center
+Mechatronic system
 Elevator
-Signals: humidity, vibration Prognosis Link
+Signals: humidity, vibration
+Prognosis Link
 110 Zenodo - TBSI Sunwoda Battery Dataset
-Shengyu, T. (2024). TBSI Sunwoda Battery Dataset [Data set]. Zenodo. https://doi.org/10.5281/zenodo.10715209 Electrical component
+Shengyu, T. (2024). TBSI Sunwoda Battery Dataset [Data set]. Zenodo. https://doi.org/10.5281/zenodo.10715209
+Electrical component
 Battery
-Signals: current, voltage Prognosis Link
+Signals: current, voltage
+Prognosis Link
+Appendix B: Classification of Publicly Available Data Sets
+[…]
+Domain Application FD D HA P Sum across domain
+Building Building 1 2
+Bridge 1
+Drive technology Aircraft engine* 5 12
+Electric motor 1 1 1
+Diesel engine 1
+Helicopter engine 1
+Propulsion system* 1 1
+Electrical component Battery 1 15 25
+Capacitor 2
+Circuit breaker 1
+Fuel cell 2
+Sensor 1 1
+Transistor 2
+Manufacturing process Planarization system 1 8
+Drilling 1
+Electrophoresis painting 1
+Milling 2 2
+Turning 1
+Material Aluminum plate 2 4
+Polymer composite 1
+Steel plate 1
+Mechanical component Anemometer 1 26
+Ball screw drive 1 1
+[…]
+Bogie 1
+Gear 1 3
+Shaft 1
+Mechatronic system Air compressor 1 8
+Air pressure system 1
+Electromechanical ball screw* 1
+Electromechanical device 1
+Elevator 1
+Hydraulic system 1
+Rock drill 1
+Validation door 1
+Process technology Filtration 5 7
+Pump 1 1
+Production system Ion mill etching tool 1 8
+Log data 1
+Production line 2 3
+Shrink-wrapper 1
+Robotic Articulated robot 4 5
+Linear robot 1
+Unknown 2 2 1 5
+all 18 35 3 54 110
+Appendix C: Signals within Publicly Available Data Sets
+Table 2: Signals used in at least two of the given data sets.
+Sorted by the sum of signal occurrences across the tasks of fault detection (FD), diagnosis (D), health assessment (HA), and prognosis (P).
+Signal FD D HA P Sum across tasks
+Vibration 4 16 0 14 34
+Current 2 9 0 22 33
+Temperature 4 8 0 21 33
+Voltage 2 4 0 19 25
+Speed 4 11 0 5 20
+Pressure 0 5 1 12 18
+Anonymized(1) 3 1 2 4 10
+Flow Rate 0 1 0 9 10
+Position 2 4 0 2 8
+Torque 2 5 0 1 8
+Operating Condition(2) 0 1 0 6 7
+Acoustic Emission 0 2 0 3 5
+Force 0 3 0 1 4
+Acceleration(3) 1 2 0 0 3
+Unknown(4) 2 0 0 1 3
+Capacity 0 0 0 2 2
+Humidity 0 1 0 1 2
+Image 1 0 0 1 2
+Impedance 0 0 0 2 2
+Inspection Data 0 2 0 0 2
+Power 1 1 0 0 2
+Strain 0 1 0 1 2
+(1): signal values changed/transformed,  (2): varying operating conditions during the life
+(3): in terms of movement, not vibration, (4): signal (or sensor) type unknown